### v1

# Quasi-perfect blowups detect regularity

Abstract.
This work gives a new characterization of regular Noetherian schemes (more generally algebraic spaces) as those for which blowups at closed points are quasi-perfect. In particular, regularity is controlled by the property that alterations or modifications of the scheme (or algebraic space) are quasi-perfect. Additionally, we study the local behavior of quasi-perfect morphisms, showing that for proper morphisms this property can be detected via stalks, and completions and (strict) Henselizations of those. Furthermore, we show that the locus of points where a proper morphism is quasi-perfect forms a Zariski open set.
Key words and phrases:
Derived category, quasi-perfect morphism, blowup, regularity, generation (for triangulated categories)
2020 Mathematics Subject Classification:
14A30 (primary), 14B05, 18G80, 13D0
1. Introduction
The zoo of singularities consists of many exotic creatures. To this day the study of its inhabitants is an active area of research. A very useful and helpful way to understand them is through geometrical and homological techniques. Typically, one associates to each inhabitant invariants. Such an invariant could be a category; it is natural to then try and study the information contained within, its deeper structure.
An important category associated to any Noetherian scheme $X$ is its derived category $D_{\operatorname{qc}}(X)$ of complexes of $\mathcal{O}_{X}$-modules with quasi-coherent cohomology. This ‘big’ category contains two important (essentially) small subcategories. Firstly, the category of perfect complexes, denoted $\operatorname{Perf}(X)$, and secondly the category of bounded complexes with coherent cohomology, denoted $D^{b}_{\operatorname{coh}}(X)$. One has the containment $\operatorname{Perf}(X)\subseteq D^{b}_{\operatorname{coh}}(X)$ which controls the regularity of the scheme: these categories coincide if and only if the scheme is regular. Furthermore, it is the latter which contains the skyscraper sheaves at closed points and it is the inclusion of these in the former, i.e. the skyscraper sheaves being perfect, which controls regularity at the corresponding points. It is this fact that leads to the main results in this work.
Probing $D^{b}_{\operatorname{coh}}(X)$ using categorical incarnations of geometric operations allows one to study its structure more precisely and uncover more information contained within. A key example is utilizing the derived pushforward along a proper morphisms $\mathbf{R}f_{\ast}\colon D^{b}_{\operatorname{coh}}(Y)\to D^{b}_{\operatorname{coh}}(X)$ to characterize certain singularities. Important examples include
  • rational singularities [Art66, Lip69] and their ‘mild’ variants [Lan25],
  • Du Bois singularities [DB81, Ste81, Ste83],
  • derived splinters [Bha12] and their ‘birational’ variants (following Kovács) [Lyu22, Lan25].
Such singularities are often considered to be ‘tame’ as they usually enjoy favorable properties; e.g. derived splinters being normal or Cohen–Macaulay in many settings. Loosely speaking, these type of singularities represent ‘slight deviations’ away from regularity in regards to (cohomological) behavior.
The story of this note began with [MS20]. There the authors showed that a local domain $R$, which is essentially of finite type over a field, is regular if and only if for every alteration $f\colon X\to\operatorname{Spec}(R)$ from a regular scheme one has $\mathbf{R}f_{\ast}\mathcal{O}_{X}\in\operatorname{Perf}(X)$. An analogue of this result in mixed characteristic was posed as a question (see [MS20, Question 1.1]). Our initial goal was to answer Ma and Schwede’s question. Whilst we were unsuccessful in that avenue, in the process it became clear that one can characterize regularity by having ‘enough probing morphisms’ be quasi-perfect in the sense of Lipman–Neeman [LN07]. Here quasi-perfect means that the derived pushforward preserves perfect complexes (see Definition 2.2).
As this characterization has, to the best of our knowledge, not appeared in the literature before, it felt worthwhile to write this short note to bring it to peoples attention. In fact, all of the results below are in the generality of algebraic spaces; but, we restrict to schemes in this introduction for ease.
Theorem A (see Theorem 3.3 ).
A Noetherian scheme $X$ is regular if, and only if, the blowup of $X$ along any closed point is a quasi-perfect morphism.
Although A does not answer [MS20, Question 1.1], it provides new and interesting characterizations of regularity: one can characterize regularity by probing with alterations or modifications.
Corollary B (see Corollary 3.6).
An integral Noetherian scheme $X$ is regular if, and only if, every alteration (resp. modification) of $X$ is quasi-perfect.
Notably, we do not require the prior existence of an alteration or modification from a regular scheme. Therefore, it offers a ‘resolution free’ characterization of regularity; however, see Remark 3.7 for how one could show regularity using the existence of a ‘quasi-perfect resolution’. As a further application of A, we give a new proof that regularity descends along fpqc covers (see Corollary 3.11).
In order to prove A we use generation techniques for triangulated categories. Such techniques have been very successful in (cohomologically) characterizing singularities; for instance, the types of singularities mentioned above [LV25, LMV25, Lan25]. The crucial observation, in our case, is that one can use the skyscraper sheaf of a closed point to detect regularity at that point (we also extend this to algebraic spaces in Lemma 3.2, which seems to not have been written down before).
As A and B show that one can use quasi-perfectness of proper morphism to detect regularity, a natural next question is whether one can ‘practically’ detect quasi-perfectness. To this end, we show that this can be checked stalk-wise on the target, that is a proper morphisms is quasi-perfect if and only if all the base changes to the local rings are quasi-perfect. In particular, checking quasi-perfectness of proper morphisms reduces to showing this for morphisms with target the spectrum of a local ring $(R,\mathfrak{m})$. Furthermore, we show that one can replace $R$ by either its $\mathfrak{m}$-adic completion $\widehat{R}$ or the (strict) Henselization $R^{h}$ (resp. $R^{sh}$); neither affects the quasi-perfectness. Thus, all in all we have the following, our last main result.
Theorem C (see Theorem 3.14).
Let $f\colon Y\to X$ be a proper morphism of schemes with $X$ separated and Noetherian. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#},\widehat{\mathcal{O}}_{X,\#},\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
  1. (1)
    $f$ is quasi-perfect
  2. (2)
    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in X$.
In order to prove this theorem, we prove some lemmas, of perhaps independent interest, that allow one to check quasi-perfectness using specific types of flat covers (see Lemmas 3.9 and 3.10). Furthermore, as a corollary we show that the locus of points $p\in X$ for which the base change to $\mathcal{O}_{X,p}$ is quasi-perfect (or equivalently base change to the completion or (strict) Hensalization of the local ring) forms a Zariski open subset of $X$, see Corollary 3.15. An interesting application (which to the best of our knowledge was previously unknown) is that the property of the regular locus containing a nonempty open subset, i.e. $X$ being JJ-0, descends along proper surjective morphisms quite generally (we have a mild reducedness condition), see Proposition 3.17 for the details.
Let us also say a word concerning the restriction to a proper morphism in C. In view of A this is mild as the morphisms appearing there are proper. In fact, for a separated morphisms of finite type between Noetherian schemes, it being quasi-perfect implies that it is proper, see e.g. [Nee21, Proposition 0.20]; so in practice it is very mild indeed. However, the real reason we need to restrict to this is that in the proofs we need to leverage certain local-to-global techniques which require working with bounded and coherent objects (e.g. [AIL10, Theorem 4.1]).
We end this introduction with a few words on the generalization to algebraic spaces. Although we stated the above results for schemes, these are proven below for algebraic spaces. This is done Section 3 after a short recollection of necessary definitions and notation in Section 2. We expect the more general results to be beneficial for future work for those studying singularities of algebraic spaces which is recently been gaining traction. For example, recent progress has been made in establishing the minimal model program for excellent algebraic spaces in characteristic zero, see [LM22].
Acknowledgments.
De Deyn was supported by ERC Consolidator Grant 101001227 (MMiMMa). Lank was supported under the ERC Advanced Grant 101095900-TriCatApp. The authors thank Leovigildo Alonso Tarrío, Fei Peng, and Karl Schwede for looking at an earlier version of this work. Lank would like to dedicate this work to Boz, whose joyous and warm memories left behind will be forever remembered.
2. Preliminaries
We make some recollections in this section, mainly to keep things self-contained and to fix notation and terminology.
2.1. Generation
Let $\mathcal{T}$ be a triangulated category. A triangulated subcategory of $\mathcal{T}$ is called thick if it is closed under direct summands (and so automatically strict, i.e. closed under isomorphisms). For any collection of objects $S$ of $\mathcal{T}$, let $\operatorname{add}(S)$ denote the smallest strictly full subcategory of $\mathcal{T}$ containing $S$ that is closed under shifts, finite coproducts and direct summands. Inductively define
$\langle S\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&n=0,\\ \operatorname{add}(S)&n=1,\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&n>1.\end{cases}$
An important feature of this construction is that $\langle S\rangle:=\cup_{n\geq 0}\langle S\rangle_{n}$ is the smallest thick triangulated subcategory of $\mathcal{T}$ containing $S$. Observe also that it is possible to rewrite the definition of $\langle\mathcal{S}\rangle_{n}$ using $\operatorname{Hom}(\langle\mathcal{S}\rangle_{1},\langle\mathcal{S}\rangle_{n-1})$ instead (so there is no choice ‘left vs. right’).
2.2. Algebraic Spaces
Our conventions follow that of the Stacks Project [Sta25]. Throughout we work with algebraic spaces in the ‘absolute’ setting, i.e. over the base scheme $\operatorname{Spec}(\mathbb{Z})$; but, of course, everything remains valid when working over an arbitrary base.
Let $X$ be an algebraic space throughout this subsection . We denote by $|X|$ the set of points of $X$, which obtains a topology by giving it the quotient topology via an étale surjective morphism from a scheme; see e.g. [Sta25, Tag 03BT]. The algebraic spaces that are considered in this work will always be decent in the sense of [Sta25, Tag 03I7]. This condition ensures that there is a good relationship between the topology of $|X|$ and properties of $X$ itself. Note that quasi-separated algebraic spaces are automatically decent; in fact, for locally Noetherian algebraic spaces being descent is equivalent to being quasi-separated [Sta25, Tag 0BB6].
Defining irreducibility for algebraic spaces is a bit subtle as this notion is not étale local. Therefore, recall that an algebraic space $X$ is called irreducible when its underlying topological spaces $|X|$ is irreducible; it is not clear whether one can characterize this using coverings from schemes (e.g. using étale surjective morphisms from schemes). An integral algebraic spaces, then, is one that is irreducible, reduced and decent [Sta25, Tag 0AD3].
Modules and derived categories
Let $X_{\acute{e}t}$ denote the small étale site of $X$. The assignment $U\mapsto\Gamma(U,\mathcal{O}_{U})$ defines a sheaf of rings on $X_{\acute{e}t}$, denoted $\mathcal{O}_{X}$ and called the structure sheaf (of $X$). Consequently, it makes sense to speak of sheaves $\mathcal{O}_{X}$-modules on $X_{\acute{e}t}$; denote the (Grothendieck) Abelian category of such sheaves by $\operatorname{Mod}(X)$. The Abelian subcategories of quasi-coherent resp. coherent $\mathcal{O}_{X}$-modules are denoted by $\operatorname{Qcoh}(X)$ resp. $\operatorname{Coh}(X)$.
Next, let $D(X):=D(\operatorname{Mod}(X))$ denote the (unbounded) derived category of $\operatorname{Mod}(X)$ and let $D_{\operatorname{qc}}(X)$ resp. $D_{\operatorname{coh}}(X)$ be the full subcategories consisting of complexes with quasi-coherent resp. coherent cohomology sheaves. Furthermore, let $D^{\flat}_{\#}(X):=D^{\flat}(X)\cap D_{\#}(X)$ for $\flat\in\{+,-,b,\geq n,\dots\}$ and $\in\{\operatorname{qc},\operatorname{coh}\}$, where the superscripts have to interpreted in the usual fashion.
Any morphisms $f\colon Y\to X$ of algebraic spaces induces a morphism of ringed topoi [Sta25, Tag 03G8], which similarly as in the case of schemes leads to a derived adjunction $\mathbf{L}f^{\ast}\colon D(X)\rightleftarrows D(Y)\colon\mathbf{R}f_{\ast}$; see e.g. [Sta25, Tag 07A6]. When $f$ is quasi-compact and quasi-separated, the derived pushforward preserves objects with quasi-coherent cohomology (see e.g. [Sta25, Tag 08FA]) and so the adjunction restricts to $D_{\operatorname{qc}}$, i.e. we have an adjunction $\mathbf{L}f^{\ast}\colon D_{\operatorname{qc}}(X)\rightleftarrows D_{\operatorname{qc}}(Y)\colon\mathbf{R}f_{\ast}$.
Perfect complexes
Perfect complexes can be defined for any ringed site [Sta25, Tag 08G4], and so, in particular, for $X$ by looking at its small étale site. Recall that a complex is strictly perfect if it is a bounded complex with each term a direct summand of a finite free module. A complex is perfect if it is locally strictly perfect. The subcategory of $D_{\operatorname{qc}}(X)$ consisting of perfect complexes is denoted by $\operatorname{Perf}(X)$; this is a thick subcategory.
When $X$ is quasi-compact and quasi-separated, the perfect complexes are exactly the compact objects of $D_{\operatorname{qc}}(X)$, see [Sta25, Tag 09M8]. Recall that an object in a triangulated category with arbitrary direct sums is said to be compact if the covariant Hom out of that object commutes with direct sums. Moreover, in this case, $D_{\operatorname{qc}}(X)$ is compactly generated by a single perfect complex, see [Sta25, Tag 0AEC]. Here $P$ compactly generating means that it is compact and any non-zero object admits a non-zero morphism from $P$, any such object is called a compact generator. It is worthwhile noting that $P$ being a compact generator for $D_{\operatorname{qc}}(X)$ is equivalent to it being a classical generator for $\operatorname{Perf}(X)$, i.e.  $\operatorname{Perf}(X)=\langle P\rangle$. This follows as $D_{\operatorname{qc}}(X)$ is compactly generated, see e.g. [Sta25, Tag 09SR].
Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g. [HR17, Lemma 4.1] or [DLMRP25, Lemma 2.2]. For the definition of the the fpqc topology, see [Sta25, Tag 03MP]; important examples of fpqc covers are fppf and so also smooth and étale covers.
Lemma 2.1.
Let $X$ be an algebraic space and $E\in D_{\operatorname{qc}}(X)$. Then $E$ is perfect if and only if there exists an fpqc cover $\{f_{i}\colon X_{i}\to X\}$ such that each $f^{\ast}_{i}E$ is perfect.
Local rings
Let $\overline{x}$ be a geometric point of $X$, i.e. $\overline{x}$ is a morphism from the spectrum of an algebraically closed field. For any presheaf $\mathcal{F}$ on $X_{\acute{e}t}$ one defines the stalk of $\mathcal{F}$ at $\overline{x}$ as
$\mathcal{F}_{\bar{x}}:=\mathop{\mathrm{colim}}\nolimits_{(U,\overline{u})}\mathcal{F}(U)$
where the colimit runs over the étale neighborhoods of $\overline{x}$ in $X$. Taking stalks at geometric points are exactly the points of the small étale site of $X$ (see e.g. [Sta25, Tag 04K6]). The étale local ring of $X$ at $\overline{x}$ is by definition the stalk of the structure sheaf at $\overline{x}$, i.e. $\mathcal{O}_{X,\overline{x}}:=(\mathcal{O}_{X})_{\overline{x}}$. This is in fact a local ring (as can be seen from the following paragraph).
To make the latter more explicit, pick an étale neighborhood $(U,\overline{u})$ of $\overline{x}$ where $U$ a scheme. Then, from [Sta25, Tag 04KF], we have isomorphisms
(2.1) $\mathcal{O}_{X,\overline{x}}\cong\mathcal{O}_{U,\overline{u}}\cong\mathcal{O}_{U,u}^{sh},$
where $\mathcal{O}_{U,u}^{sh}$ is the strict Henselization of $\mathcal{O}_{U,u}$ and $u\in U$ is the point at which $\overline{u}$ is centered. We refer to [Sta25, Tag 0BSK] for the definition of (strict) Henselizations of local rings. Because of Equation 2.1, for arbitrary $x\in|X|$, by abuse of notation we sometimes denote $\mathcal{O}_{X,x}^{sh}:=\mathcal{O}_{X,\overline{x}}$ where $\overline{x}$ is any geometric point lying over $x$.
Lastly, let us briefly define the Henselian local ring at a point $x\in|X|$. Let
$\mathcal{O}_{X,x}^{h}:=\mathop{\mathrm{colim}}\nolimits_{(U,u)}\Gamma(U,\mathcal{O}_{U})$
where now the colimit is taken over the ‘elementary étale neighborhoods’ (see [Sta25, Tag 0BGU] for the definition). As in the previous paragraph one can make this more explicit in terms of a covering. Indeed, by [Sta25, Tag 0EMY], picking an elementary etale neighborhood $(U,u)\to(X,x)$ one has $\mathcal{O}_{X,x}^{h}\cong\mathcal{O}_{U,u}^{h}$, where $\mathcal{O}_{U,u}^{h}$ is the Henselization of the local ring $\mathcal{O}_{U,u}$ of $U$ at $u$. In particular,
(2.2) $(\mathcal{O}_{X,x}^{h})^{sh}=\mathcal{O}_{X,x}^{sh},$
see e.g. [Sta25, Tag 0EMZ].
Residue fields
Pick a point $x\in|X|$. When $X$ is decent there exists, by [Sta25, Tag 03K4], a unique monomorphism $\operatorname{Spec}(k)\to X$ representing $x$. The field $k$ appearing is called the residue field of $X$ at $x$ and is denoted $\kappa(x)$.
Regularity
Recall that $X$ is regular if there exists an étale surjective morphism from a regular scheme, see e.g. [Sta25, Tag 03E5]. Furthermore, a point $x\in|X|$ is regular if there exists an étale morphism $f\colon U\to X$ from a scheme and a point $u\in U$ such that $f(u)=x$ and the stalk $\mathcal{O}_{U,u}$ is a regular local ring; this is well-defined by [Sta25, Tag 0AH7]. It is straightforward to see that an algebraic space is regular if and only if all of its points are regular. Moreover, by [Sta25, Tag 0AHA], a point $x\in|X|$ is regular if and only if the étale local ring $\mathcal{O}_{X,\overline{x}}$ is a regular local ring for any (equivalently some) geometric point $\overline{x}$ lying over $x$.
When $X$ is a Noetherian algebraic space, it follows from [DLMRP25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$. Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLMRP25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e. $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.
2.3. Morphisms
We say a morphism of integral algebraic spaces $f\colon Y\to X$ is
  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [Sta25, Tag 0ACU] for some characterizations and
  • an alteration if it is proper and generically finite, see e.g. [Sta25, Tag 0AD7] for some characterizations.
Usually, in this work the target (and hence, source) of the morphism will be Noetherian. Note that, for either type, the morphism above is, in fact, surjective. This follows as they are by definition dominant (this is implicit in generically finite/an isomorphism) and a dominant proper morphism is surjective. The modifications we will be interested in below arise as blowups. See [Sta25, Tag 085P] some background on blowups of algebraic spaces.
Let $f\colon Y\to X$ now be more generally any morphism of quasi-compact quasi-separated algebraic spaces. Extending [LN07] to algebraic spaces we obtain the following.
Definition 2.2.
We say $f$ is quasi-perfect if $\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$, i.e. pushforward preserves compacts.
The pushforward preserving compact objects has many different characterizations. Let us just mention [BDS16, Theorem 3.3] for a nice list of characterizations when viewing the above through a tensor triangular lens.
Warning.
A morphism of schemes being quasi-perfect should not be confused with it being a ‘perfect morphisms’, another notion that appears in the literature. These notions coincide, for example, for a proper morphisms of Noetherian schemes [LN07, Theorem 1.2], but differ in general.
3. Results
To start, we note an ‘obvious’ lemma which is recorded for convenience.
Lemma 3.1.
Let $f\colon Y\to X$ be a morphisms of quasi-compact quasi-separated algebraic spaces with $Y$ nonempty. Then $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is nonzero; in particular, $\mathbf{R}f_{\ast}G$ is nonzero for any compact generator $G$ of $D_{\operatorname{qc}}(Y)$.
Proof.
The second claim follows from the first as $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is finitely built from $\mathbf{R}f_{\ast}G$ (i.e. $\mathcal{O}_{Y}\in\langle G\rangle$ and so also after applying $\mathbf{R}f_{\ast}$). To see the first claim, observe that it follows from pullback/pushforward adjunction that
$0$ $\not=H^{0}(Y,\mathcal{O}_{Y})$
$\cong\operatorname{Ext}^{0}(\mathcal{O}_{Y},\mathcal{O}_{Y})$
$\cong\operatorname{Ext}^{0}(\mathbf{L}f^{\ast}\mathcal{O}_{X},\mathcal{O}_{Y})$
$\cong\operatorname{Ext}^{0}(\mathcal{O}_{X},\mathbf{R}f_{\ast}\mathcal{O}_{Y})$
and so $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is nonzero. ∎
The next should be compared to [Sta25, Tag 00OC] and gives a ‘spacy’ version of this.
Lemma 3.2.
Let $X$ be a decent (e.g. a Noetherian) algebraic space. Suppose $x\in|X|$ is a closed point such that $i_{x,*}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}$ is perfect, where $i_{x}\colon{\operatorname{Spec}}(\kappa(x))\to X$ denotes the residue field of $X$ at $x$. Then $x$ is a regular point of $X$.
Proof.
Let $(U,u)\to(X,x)$ be the ‘elementary étale neighborhood’ of [Sta25, Tag 0BBP], i.e. $U$ is an affine scheme, $u$ is the only point of $U$ lying over $x$, and the induced homomorphism of residue fields $\kappa(x)\to\kappa(u)$ is an isomorphism. We claim that the pullback of $x$ under $f\colon U\to X$ can be identified with $u$, i.e. we have a cartesian square
[…]
Indeed, let $U_{x}$ denote the pullback. It follows from categorical nonsense that $U_{x}\to U$ is a monomorphism (as $\operatorname{Spec}(\kappa(x))\to X$ is). Moreover, as $U_{x}\to\operatorname{Spec}\kappa(x)$ is étale, it is a disjoint union of fields. However, as each field corresponds to a distinct point lying over $x$ (they all go to different points as monomorphisms of schemes are injective), $U_{x}$ is the spectrum of a field. Consequently, we must have $U_{x}=\operatorname{Spec}(\kappa(u))$ as $U_{x}\to U$ is a monomorphism. In particular, it also follows that $u$ is necessarily a closed point as $i_{x}$ is a closed immersion by e.g. [Sta25, Tag 0H1R & Tag 0H1U].
Next, using e.g. flat base change, one sees
$f^{*}i_{x,*}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}=i_{u,*}\mathcal{O}_{\operatorname{Spec(\kappa(u))}}$
and so also perfect. Taking the stalk at $u$, one has the residue field $\kappa(u)$ is a perfect $\mathcal{O}_{U,u}$-module. Thus, $u$ is a regular point by [Sta25, Tag 00OC] and so $x$ is a regular point of $X$ by definition. ∎
Theorem 3.3.
A Noetherian algebraic space $X$ is regular if, and only if, the blowup of $X$ along any closed point is a quasi-perfect morphism.
Proof.
If $X$ is regular, then for any proper $f$ we have
$\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\subseteq D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X).$
Hence, any blowup is quasi-perfect as these are proper.
For the converse, it suffices to show that any closed point is regular (see e.g. [DLMRP25, Lemma 3.2]). So, let $p\in|X|$ be a closed point and denote by $i\colon\operatorname{Spec}(\kappa(p))\to X$ its associated residue field. As the point is closed, $i$ is a closed immersion; see e.g. [Sta25, Tag 0H1R & Tag 0H1U]. Consider the fibered square
[…]
where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor. The hypothesis tells us that $f$ is a quasi-perfect morphism.
As the exceptional divisor $E$ is an effective Cartier divisor on $X^{\prime}$, we have a short exact sequence
$0\to\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to 0$
from which we obtain a distinguished triangle in $D^{b}_{\operatorname{coh}}(X^{\prime})$,
(3.2) $\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to\mathcal{O}_{X^{\prime}}(-E)[1].$
Furthermore, as $\mathcal{O}_{X^{\prime}}(-E)$ is an invertible sheaf, and so perfect, we have $j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X^{\prime})$.
Hence, as $f$ is quasi-perfect by assumption, we know that $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$. Moreover, from the commutativity of Equation 3.1 and Lemma 3.1, we see that
$0$ $\not=\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$
$\cong i_{\ast}\mathbf{R}g_{\ast}\mathcal{O}_{E}$
$\cong i_{\ast}(\text{sum of shifts of }\mathcal{O}_{\operatorname{Spec}(\kappa(p))})$
$\cong\text{sum of shifts of }i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$
It follows that $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect. Hence $p$ is regular by Lemma 3.2 and we are done. ∎
Remark 3.4.
It is worthwhile to stress the following: the blowup at a closed point being quasi-perfect implies that that point is regular. Indeed, this is essentially the content of the above proof.
Lemma 3.5.
The blowup $f\colon X^{\prime}\to X$ of an integral algebraic space along a nonzero quasi-coherent ideal sheaf is integral. In particular, when the ideal is of finite type, it is a modification.
Proof.
The blowup is reduced, see e.g. [Sta25, Tag 085W], as the latter is an étale local property. Moreover, since blowups are representable, it follows that $X^{\prime}$ is decent (see e.g. [Sta25, Tag 0ABT]).
Thus, it remains to show that $|X^{\prime}|$ is irreducible, which is more subtle as this notion is not étale local. Let $Z$ be the center of the blow-up. We have that $f^{-1}(Z)$ is an effective Cartier divisor on $X^{\prime}$ and that $f$ induces an isomorphism $f^{-1}(X\setminus Z)=X\setminus Z$; see e.g. [Sta25, Tag 085T]. When $|X|$ is irreducible so is the open subset $|X\setminus Z|$ and so $|f^{-1}(X\setminus Z)|=|X^{\prime}\setminus f^{-1}(Z)|$ is also. However, $f^{-1}(Z)$ being effective Cartier implies that the open subspace $X^{\prime}\setminus f^{-1}(Z)$ is scheme-theoretically dense in $X^{\prime}$, see [Sta25, Tag 083S]. In particular, the closure of $|X^{\prime}\setminus f^{-1}(Z)|$ (in $|X^{\prime}|$) is $|X^{\prime}|$ and so the latter is also irreducible.
To see that $f$ is a modification when the ideal is of finite type, note that in this case $f$ is proper by [Sta25, Tag 085Z]. ∎
Corollary 3.6.
An integral Noetherian algebraic space $X$ is regular if, and only if, every alteration (resp. modification) of $X$ is quasi-perfect.
Proof.
This follows immediately from Theorem 3.3. Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma 3.5. Moreover, as modifications are alterations the case of alterations also follows. ∎
Remark 3.7.
In the special case where there exists an alteration or modification $f\colon Y\to X$ of algebraic spaces where $Y$ is regular and $f$ quasi-perfect, then one can show $X$ regular more easily using ‘proper descent’. For example, this follows from [HLLP25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity. So, the novelty of Corollary 3.6 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
Next come some lemmas that allow one to check quasi-perfectness using flat covers.
Lemma 3.8.
A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if, and only if, $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
Proof.
If $f$ is quasi-perfect, it is clear that $\mathbf{R}f_{\ast}P\in\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$. For, the converse direction, let $P$ be compact generator on $Y$. Our hypothesis is $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$. The desired claim then follows from the following string of inclusions:
$\mathbf{R}f_{\ast}\operatorname{Perf}(Y)=\mathbf{R}f_{\ast}\langle P\rangle\subseteq\langle\mathbf{R}f_{\ast}P\rangle\subseteq\operatorname{Perf}(X).\qed$
Lemma 3.9.
Let $f\colon Y\to X$ be a morphism of quasi-compact quasi-separated algebraic spaces and let $s\colon U\to X$ be a flat morphism (from an algebraic space). Consider the fibered square
[…]
  1. (1)
    If $f$ is quasi-perfect and $s$ is affine, then $f^{\prime}$ is quasi-perfect.
  2. (2)
    If $s$ is surjective and $f^{\prime}$ is quasi-perfect, then $f$ is quasi-perfect.
Proof.
To show the first claim, first note that when $s$ is an affine morphism so is its base change $s^{\prime}$. Consequently, $\mathbf{R}s^{\prime}_{\ast}=s^{\prime}_{\ast}$ is conservative (this statement is local on the target and so reduces to the case of schemes) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g. [Sta25, Tag 0E4R]). Thus, as $f$ is assumed quasi-perfect, $\mathbf{R}f_{\ast}Q$ is perfect and so also is its pullback $s^{\ast}\mathbf{R}f_{\ast}Q$. From flat base change, it follows that
(3.4) $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q\cong s^{\ast}\mathbf{R}f_{\ast}Q.$
Hence, Lemma 3.8 implies $f^{\prime}$ is a quasi-perfect morphism.
For the second claim, assume now that $f$ is surjective and $f^{\prime}$ is quasi-perfect. Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation 3.4. As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact. Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma 2.1. By Lemma 3.8, we have that $f$ is quasi-perfect, which completes the proof. ∎
The following is a variation on the previous.
Lemma 3.10.
Let $f\colon Y\to X$ be a morphism of quasi-compact quasi-separated algebraic spaces with $X$ an affine scheme. Suppose $s\colon U\to X$ is a flat surjective morphism from an affine scheme. Then $f$ is quasi-perfect if and only if the base change of $f$ along $s$ is quasi-perfect.
Proof.
The forward direction follows from Lemma 3.9. For the converse, denote the base change of $f$ along $s$ as in Equation 3.3. Again, choose a compact generator $Q$ of $Y$. By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation 3.4, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect. However, as $s$ corresponds to a faithfully flat ring map, it follows from [Sta25, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect. Thus, Lemma 3.8 shows that $f$ is quasi-perfect. ∎
Using Lemma 3.9, we obtain an alternative proof of the fact that regularity can be checked fpqc locally.
Corollary 3.11.
Let $X$ be a separated Noetherian algebraic space. If there is an fpqc cover $\{f_{i}\colon U_{i}\to X\}_{i\in I}$ where each $U_{i}$ is regular, then $X$ is regular.
Proof.
First, as $X$ is quasi-compact, we can reduce to an fpqc cover with finitely many components (i.e. $|I|<\infty$). Secondly, if needed, we can refine the cover so that each $U_{i}$ is an affine scheme (e.g. take étale surjective morphisms from affine schemes). Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is affine by [Sta25, Tag 08GB] as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme. To show $X$ is regular, it suffices by Theorem 3.3 to show that blowups along nonzero ideal sheaves are quasi-perfect. However, as blowups commute with flat base change, see e.g. [Sta25, Tag 0805], we have that the base change to $U$ is again such a blow-up. Consequently, as $U$ is regular, it is quasi-perfect. Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma 3.9, showing the claim. ∎
Remark 3.12.
In fact, when looking at fppf covers it is enough to only assume that $X$ is quasi-compact and decent in the previous by making use of some descent. Indeed, it follows from the existence of an fppf cover as in the statement that $X$ is locally Noetherian (as this is fppf local and regular algebraic spaces are always locally Noetherian) and hence Noetherian.
There are three important cases of faithfully flat morphisms that we will apply Lemma 3.10 to.
Example 3.13.
Let $(R,\mathfrak{m})$ be a Noetherian local ring. To this one can associate a few other Noetherian local rings:
  1. (1)
    the $\mathfrak{m}$-adic completion of $R$, denoted $\widehat{R}$,
  2. (2)
    the Henselization of $R$, denoted $R^{h}$, and
  3. (3)
    the strict Henselization of $R$, denoted $R^{sh}$.
There are local ring morphisms $R\to\widehat{R}$ and $R\to R^{h}\to R^{sh}$ which are flat and consequently faithfully flat; see e.g. [Sta25, Tag 00MC and Tag 07QM]. (As an aside, there is also a flat local ring morphism $R^{h}\to\widehat{R}$, but we do not need it).
We now move to our next main result.
Theorem 3.14.
Let $f\colon Y\to X$ be a proper morphism of algebraic spaces with $X$ separated and Noetherian. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
  1. (1)
    $f$ is quasi-perfect,
  2. (2)
    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in|X|$.
Furthermore, if $s\colon U\to X$ is an étale surjective morphism from a quasi-compact scheme, one can take $A_{\#}\in\{\mathcal{O}_{U,\#},\widehat{\mathcal{O}}_{U,\#}\}$ and let $p$ run over all points in $U$.
Proof.
[…]
($\star$) If the base change of $f$ to $\mathcal{O}_{U,p}$ is quasi-perfect for all $p\in U$, then $f$ is quasi-perfect.
The proof of ($\star$3) is similar to that of Lemma 3.9; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work. Pick a point $p\in U$ and look at the fibered square
[…]
where $s_{p}$ is the natural morphism and $f_{p}$ and $t_{p}$ are the projections.
Again, choose any compact generator $Q$ of $Y$. The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect. As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper. Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma 2.1 and $f$ is quasi-perfect by Lemma 3.8. ∎
Corollary 3.15.
Let $f\colon Y\to X$ be a proper morphism of algebraic spaces with $X$ separated and Noetherian and let $A_{\#}$ and $s\colon U\to X$ be as in Theorem 3.14. Then the locus
(3.5) $S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
Proof.
By the proof of Theorem 3.14 (i.e. using Lemma 3.10) we know that
$S_{X}:=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
$=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{h}\textrm{ is quasi-perfect}\right\}$
and
$S_{U}:=$ $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
$=$ $\left\{u\in U\mid\text{base change to }\widehat{\mathcal{O}}_{U,u}\textrm{ is quasi-perfect}\right\}.$
Let us start by showing that
(3.6) $s^{-1}(S_{X})=S_{U}.$
To this end, observe by Equation 2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$. As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma 3.10. Therefore,
$s^{-1}(S_{X})$ $=\left\{u\in U\mid\text{base change to }\mathcal{O}_{X,s(u)}^{sh}\textrm{ is quasi-perfect}\right\}$
$=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}^{sh}\textrm{ is quasi-perfect}\right\}$
$=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}=S_{U}.$
As a consequence of Equation 3.6 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open. Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
Therefore, we may and will simply assume that $X$ is an affine scheme from now on. Choose any compact generator $Q$ of $Y$. As $f$ is proper, $\mathbf{R}f_{\ast}Q$ is bounded coherent and so by [Let21, Proposition 3.5] it follows that
$S^{\prime}:=\left\{x\in X\mid(\mathbf{R}f_{\ast}Q)_{x}\in\langle\mathcal{O}_{X,x}\rangle=\operatorname{Perf}(\mathcal{O}_{X,x})\right\}$
is a Zariski open subset of $X$. Considering the fibered square
[…]
we see that, using flatness of $s_{x}$ and that it is affine so that $t_{x}^{\ast}Q$ is a compact generator for $Y\times_{X}\operatorname{Spec}(\mathcal{O}_{X,x})$,
$x\in S\iff$ $\text{base change to }\mathcal{O}_{X,x}\textrm{ is quasi-perfect}$
$\iff$ $\mathbf{R}(f_{x})_{\ast}t_{x}^{\ast}Q\text{ is perfect}$
$\iff$ $(\mathbf{R}f_{\ast}Q)_{x}=s_{x}^{\ast}\mathbf{R}f_{\ast}Q\text{ is perfect}\iff x\in S^{\prime}$
finishing the proof. ∎
Remark 3.16.
One can think of the subset $S$ appearing in Corollary 3.15 as the ‘quasi-perfect locus of $f$’. It follows from the first part of the proof that in general, i.e. regardless of properness of $f$, one can compute it étale locally (Equation 3.6) and that the specific choice of $A_{\#}$ does not matter. Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
It is also worthwhile to point out that, quite generally for any morphisms $f\colon Y\to X$ (of quasi-compact quasi-separated algebraic spaces) with separated target, we have an inclusion of the regular locus $\operatorname{reg}(X)\subseteq S$. Indeed, if $p\in|X|$ is a regular point, its étale local ring is regular and so the base change to it is quasi-perfect.
As an application of the ‘quasi-perfect locus’, we show it can be used to descent properties of the regular locus along proper morphisms. Moreover, note that the result requires no separations assumptions on the target and that the assumption on $X$ in the statement is a necessary condition due to the conclusion of the proposition.
Proposition 3.17.
Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [Sta25, Tag 04NA] for the relevant definition. whose étale local ring is reduced (e.g. this is the case when $X$ is reduced). Suppose further that $f\colon Y\to X$ is a proper surjective morphism. Then the regular locus of $X$ contains a nonempty open subset (i.e. is JJ-0) whenever $Y$ has this property.
Proof.
The proof starts by a series of reductions. Let us denote by $\eta$ the codimension zero point whose étale local ring is reduced.
  • As $X$ admits an open dense subscheme, which in particular contains $\eta$, see e.g. [Sta25, Tag 06NH], we may assume that $X$ is a scheme. Moreover, by choosing any affine open neighborhood $U$ of $\eta$ and by restricting to $f^{-1}(U)\to U$, we may assume $X$ is affine and, in particular, separated.
  • Next, let $S$ denote the open subset from Corollary 3.15; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty. It follows from Theorem 3.14 that the restriction $f^{-1}(S)\to S$ is quasi-perfect. So we may assume $f$ is quasi-perfect.
  • Let $V$ be a non-empty open subset of $Y$ contained in the regular locus, which exists by assumption. If $V\neq Y$, then $Y\setminus V$ is a properly contained closed subset of $Y$. As $f$ is proper, $f(Y\setminus V)$ is closed in $X$. Moreover, $f^{-1}(X\setminus f(Y\setminus V))\subset V$; so restricting to the open subset $X\setminus f(Y\setminus V)\subset X$ allows us to reduce to the setting where $f$ is a quasi-perfect proper surjective morphism from a regular scheme. (Note that whether or not $\eta\in X\setminus f(Y\setminus V)$ does not matter, the existence of $\eta$ was needed in the previous bullet.)
In order to finish the proof, is suffices to show that $X$ is regular. This follows from ‘proper descent’, as already indicated in Remark 3.7; but we give a self-contained proof, more in line with that of Theorem 3.3, for the convenience of the reader.
So, let $p\in X$ be a closed point and denote the associated closed immersion by $i\colon\mathrm{Spec}(\kappa(p))\to X$. Consider the fibered square
[…]
Clearly, $\kappa(p)$ being a field, $\mathbf{R}f^{\prime}_{\ast}\mathcal{O}_{Y\times_{X}}\operatorname{Spec}(\kappa(p))$ is a (non-zero, as $f$ is surjective!) finite direct sum of shifts of $\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$. Thus, commutativity of the diagram shows we have $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}\in\langle\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{1}$. Regularity of $Y$ ensures that $\operatorname{Perf}(Y)=D^{b}_{\operatorname{coh}}(Y)$ and quasi-perfectness of $f$ tells us $\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$. Consequently, $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}\in\operatorname{Perf}(X)$, which ensures that $p$ is a regular point of $X$. As we have shown every closed point of $X$ is regular, it follows that $X$ is regular as desired. ∎

### v2

@@
-Quasi-perfect blowups detect regularity
+A note on quasi-perfect morphisms
 Abstract.
-This work gives a new characterization of regular Noetherian schemes (more generally algebraic spaces) as those for which blowups at closed points are quasi-perfect.
-In particular, regularity is controlled by the property that alterations or modifications of the scheme (or algebraic space) are quasi-perfect.
-Additionally, we study the local behavior of quasi-perfect morphisms, showing that for proper morphisms this property can be detected via stalks, and completions and (strict) Henselizations of those.
-Furthermore, we show that the locus of points where a proper morphism is quasi-perfect forms a Zariski open set.
+This note records two results concerning quasi-perfect morphisms between Noetherian algebraic spaces.
+First, we give a new characterization of regular Noetherian algebraic spaces as those for which blowups at closed points are quasi-perfect.
+Secondly, we study the local behavior of quasi-perfect proper morphisms.
+In particular, we show that quasi-perfectness of a proper morphism can be detected at their étale local rings, completions, and (strict) Henselizations.
+As a corollary, the locus of points where a proper morphism is quasi-perfect is Zariski open.
 Key words and phrases:
@@
 2020 Mathematics Subject Classification:
-14A30 (primary), 14B05, 18G80, 13D0
+14A30 (primary), 14B05, 18G80, 13D09
 1. Introduction
-The zoo of singularities consists of many exotic creatures.
-To this day the study of its inhabitants is an active area of research.
-A very useful and helpful way to understand them is through geometrical and homological techniques.
-Typically, one associates to each inhabitant invariants.
-Such an invariant could be a category; it is natural to then try and study the information contained within, its deeper structure.
-An important category associated to any Noetherian scheme $X$ is its derived category $D_{\operatorname{qc}}(X)$ of complexes of $\mathcal{O}_{X}$-modules with quasi-coherent cohomology.
-This ‘big’ category contains two important (essentially) small subcategories.
-Firstly, the category of perfect complexes, denoted $\operatorname{Perf}(X)$, and secondly the category of bounded complexes with coherent cohomology, denoted $D^{b}_{\operatorname{coh}}(X)$.
-One has the containment $\operatorname{Perf}(X)\subseteq D^{b}_{\operatorname{coh}}(X)$ which controls the regularity of the scheme: these categories coincide if and only if the scheme is regular.
-Furthermore, it is the latter which contains the skyscraper sheaves at closed points and it is the inclusion of these in the former, i.e.
-the skyscraper sheaves being perfect, which controls regularity at the corresponding points.
-It is this fact that leads to the main results in this work.
-Probing $D^{b}_{\operatorname{coh}}(X)$ using categorical incarnations of geometric operations allows one to study its structure more precisely and uncover more information contained within.
-A key example is utilizing the derived pushforward along a proper morphisms $\mathbf{R}f_{\ast}\colon D^{b}_{\operatorname{coh}}(Y)\to D^{b}_{\operatorname{coh}}(X)$ to characterize certain singularities.
-Important examples include
-  • rational singularities [Art66, Lip69] and their ‘mild’ variants [Lan25],
-  • Du Bois singularities [DB81, Ste81, Ste83],
-  • derived splinters [Bha12] and their ‘birational’ variants (following Kovács) [Lyu22, Lan25].
-Such singularities are often considered to be ‘tame’ as they usually enjoy favorable properties; e.g.
-derived splinters being normal or Cohen–Macaulay in many settings.
-Loosely speaking, these type of singularities represent ‘slight deviations’ away from regularity in regards to (cohomological) behavior.
-The story of this note began with [MS20].
-There the authors showed that a local domain $R$, which is essentially of finite type over a field, is regular if and only if for every alteration $f\colon X\to\operatorname{Spec}(R)$ from a regular scheme one has $\mathbf{R}f_{\ast}\mathcal{O}_{X}\in\operatorname{Perf}(X)$.
-An analogue of this result in mixed characteristic was posed as a question (see [MS20, Question 1.1]).
-Our initial goal was to answer Ma and Schwede’s question.
-Whilst we were unsuccessful in that avenue, in the process it became clear that one can characterize regularity by having ‘enough probing morphisms’ be quasi-perfect in the sense of Lipman–Neeman [LN07].
-Here quasi-perfect means that the derived pushforward preserves perfect complexes (see Definition 2.2).
-As this characterization has, to the best of our knowledge, not appeared in the literature before, it felt worthwhile to write this short note to bring it to peoples attention.
-In fact, all of the results below are in the generality of algebraic spaces; but, we restrict to schemes in this introduction for ease.
-Theorem A (see Theorem 3.3 ).
-A Noetherian scheme $X$ is regular if, and only if, the blowup of $X$ along any closed point is a quasi-perfect morphism.
-Although A does not answer [MS20, Question 1.1], it provides new and interesting characterizations of regularity: one can characterize regularity by probing with alterations or modifications.
-Corollary B (see Corollary 3.6).
-An integral Noetherian scheme $X$ is regular if, and only if, every alteration (resp. modification) of $X$ is quasi-perfect.
-Notably, we do not require the prior existence of an alteration or modification from a regular scheme.
-Therefore, it offers a ‘resolution free’ characterization of regularity; however, see Remark 3.7 for how one could show regularity using the existence of a ‘quasi-perfect resolution’.
-As a further application of A, we give a new proof that regularity descends along fpqc covers (see Corollary 3.11).
-In order to prove A we use generation techniques for triangulated categories.
-Such techniques have been very successful in (cohomologically) characterizing singularities; for instance, the types of singularities mentioned above [LV25, LMV25, Lan25].
-The crucial observation, in our case, is that one can use the skyscraper sheaf of a closed point to detect regularity at that point (we also extend this to algebraic spaces in Lemma 3.2, which seems to not have been written down before).
-As A and B show that one can use quasi-perfectness of proper morphism to detect regularity, a natural next question is whether one can ‘practically’ detect quasi-perfectness.
-To this end, we show that this can be checked stalk-wise on the target, that is a proper morphisms is quasi-perfect if and only if all the base changes to the local rings are quasi-perfect.
-In particular, checking quasi-perfectness of proper morphisms reduces to showing this for morphisms with target the spectrum of a local ring $(R,\mathfrak{m})$.
-Furthermore, we show that one can replace $R$ by either its $\mathfrak{m}$-adic completion $\widehat{R}$ or the (strict) Henselization $R^{h}$ (resp.
-$R^{sh}$); neither affects the quasi-perfectness.
-Thus, all in all we have the following, our last main result.
-Theorem C (see Theorem 3.14).
-Let $f\colon Y\to X$ be a proper morphism of schemes with $X$ separated and Noetherian.
-Suppose $A_{\#}\in\{\mathcal{O}_{X,\#},\widehat{\mathcal{O}}_{X,\#},\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$.
-Then the following are equivalent:
+Given a quasi-compact quasi-separated morphism $f\colon Y\to X$ of schemes, Neeman’s Brown Representability Theorem provides a right adjoint $f^{\times}$ to the derived pushforward $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$.
+While $f^{\times}$ preserves small coproducts for complexes that are uniformly homologically bounded below, this property fails in general.
+Motivated by [VER69, Lemma 1 & Corollary 2], Lipman–Neeman [LN07] introduced the notion of quasi-perfect morphisms for quasi-compact quasi-separated schemes to remedy this.
+The defining property is that $f^{\times}$ preserves small coproducts on all of $D_{\operatorname{qc}}$ or, equivalently, that $\mathbf{R}f_{\ast}$ preserve perfect complexes; see [LN07, Theorem 1.2 & Proposition 2.1].
+Whilst there are numerous examples of quasi-perfect morphisms—for instance, any proper morphisms of finite Tor dimension between Noetherian schemes—quasi-perfectness generally fails.
+For example, this is the case for closed immersions into a non-regular Noetherian scheme.
+More recent developments have recast these ideas in a categorical framework.
+For example, Balmer–Dell’Ambrogio–Sanders develop a duality theory for rigidly-compactly generated tensor-triangulated categories [BDS16, Theorem 1.7].
+However, when interpreted in the setting of schemes, this perspective is largely global in nature.
+In particular, the local behavior of quasi-perfectness remains less understood; for instance, how quasi-perfectness at the local rings of the target influences quasi-perfectness globally.
+In this short note we make two novel observations regarding quasi-perfect morphisms between algebraic spaces. The first observation is how one can use these morphisms to characterize regularity.
+Proposition (Proposition˜3.3).
+A Noetherian algebraic space $X$ is regular if and only if the blowup of $X$ along any closed point is a quasi-perfect morphism.
+To the best of our knowledge, this characterization has not appeared before in the literature even for schemes.
+The simple, but crucial, observation is that a closed point on a Noetherian algebraic space is regular if and only if the blowup at that point is quasi-perfect.
+This follows from the fact that perfectness of the skyscraper sheaf at a closed point detects regularity of that point; we prove this in Lemma˜3.1, which seems to not have been written down for algebraic spaces before and might be of independent interest.
+Furthermore, as an application of Proposition˜3.3, we give a new proof that regularity descends along fpqc covers (see Corollary˜4.5).
+The second observation concerns the local behavior of quasi-perfectness for proper morphisms. This is summarized in the following result.
+Theorem (Theorem˜4.8).
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
   1. (1)
-    $f$ is quasi-perfect
+    $f$ is quasi-perfect,
   2. (2)
-    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in X$.
-In order to prove this theorem, we prove some lemmas, of perhaps independent interest, that allow one to check quasi-perfectness using specific types of flat covers (see Lemmas 3.9 and 3.10).
-Furthermore, as a corollary we show that the locus of points $p\in X$ for which the base change to $\mathcal{O}_{X,p}$ is quasi-perfect (or equivalently base change to the completion or (strict) Hensalization of the local ring) forms a Zariski open subset of $X$, see Corollary 3.15.
-An interesting application (which to the best of our knowledge was previously unknown) is that the property of the regular locus containing a nonempty open subset, i.e.
-$X$ being JJ-0, descends along proper surjective morphisms quite generally (we have a mild reducedness condition), see Proposition 3.17 for the details.
-Let us also say a word concerning the restriction to a proper morphism in C.
-In view of A this is mild as the morphisms appearing there are proper.
-In fact, for a separated morphisms of finite type between Noetherian schemes, it being quasi-perfect implies that it is proper, see e.g.
-[Nee21, Proposition 0.20]; so in practice it is very mild indeed.
-However, the real reason we need to restrict to this is that in the proofs we need to leverage certain local-to-global techniques which require working with bounded and coherent objects (e.g.
+    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in|X|$.
+Furthermore, if $s\colon U\to X$ is an étale surjective morphism from a quasi-compact scheme, one can take $A_{\#}\in\{\mathcal{O}_{U,\#},\widehat{\mathcal{O}}_{U,\#}\}$ and let $p$ run over all points in $U$.
+Theorem˜4.8 provides a mechanism for detecting quasi-perfectness by reducing the question to targets that are local rings.
+The proof relies on several auxiliary lemmas, potentially of independent interest, which allow quasi-perfectness to be checked using specific classes of flat covers (see Lemmas˜4.2 and 4.3).
+As a corollary, we show that the locus of points $p$, for which the base change to one of the classes of local rings $A_{p}$ from Theorem˜4.8 is quasi-perfect, forms a Zariski open subset.
+Corollary (Corollary˜4.9).
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
+$S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
+is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
+One can think of $S$ as the ‘the quasi-perfect locus of $f$’.
+The fact it is always open for proper morphisms appears to be new even for schemes.
+As an application, we show that the property that the regular locus contains a nonempty open subset (i.e.
+that $X$ is JJ-0) descends along proper surjective morphisms under mild hypotheses, which are satisfied if, for example, the target is reduced (see Proposition˜4.11).
+Finally, we comment on the restriction to proper morphisms in Theorem˜4.8.
+In view of Proposition˜3.3, this is mild since the morphisms appearing there are proper.
+In fact, for a separated morphism of finite type between Noetherian schemes, quasi-perfectness already implies properness (see [NEE21, Proposition 0.20]); so this is quite a mild restriction.
+The properness assumption is required in our arguments to employ local-to-global techniques involving bounded coherent objects (e.g.
 [AIL10, Theorem 4.1]).
-We end this introduction with a few words on the generalization to algebraic spaces.
-Although we stated the above results for schemes, these are proven below for algebraic spaces.
-This is done Section 3 after a short recollection of necessary definitions and notation in Section 2.
-We expect the more general results to be beneficial for future work for those studying singularities of algebraic spaces which is recently been gaining traction.
-For example, recent progress has been made in establishing the minimal model program for excellent algebraic spaces in characteristic zero, see [LM22].
+We expect the more general results to be useful in future work on singularities of algebraic spaces, a topic that has recently been gaining traction.
+For example, recently the minimal model program has been established for excellent algebraic spaces in characteristic zero [LM22].
 Acknowledgments.
@@
 Let $\mathcal{T}$ be a triangulated category.
+Recall that a strictly full subcategory is a full subcategory closed under isomorphisms.
 A triangulated subcategory of $\mathcal{T}$ is called thick if it is closed under direct summands (and so automatically strict, i.e.
@@
 For any collection of objects $S$ of $\mathcal{T}$, let $\operatorname{add}(S)$ denote the smallest strictly full subcategory of $\mathcal{T}$ containing $S$ that is closed under shifts, finite coproducts and direct summands.
-Inductively define
+Following [BV03], inductively, define
 $\langle S\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&n=0,\\ \operatorname{add}(S)&n=1,\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&n>1.\end{cases}$
 An important feature of this construction is that $\langle S\rangle:=\cup_{n\geq 0}\langle S\rangle_{n}$ is the smallest thick triangulated subcategory of $\mathcal{T}$ containing $S$.
-Observe also that it is possible to rewrite the definition of $\langle\mathcal{S}\rangle_{n}$ using $\operatorname{Hom}(\langle\mathcal{S}\rangle_{1},\langle\mathcal{S}\rangle_{n-1})$ instead (so there is no choice ‘left vs.
-right’).
+Observe also that it is possible to rewrite the definition of $\langle\mathcal{S}\rangle_{n}$ using $\operatorname{Hom}(\langle\mathcal{S}\rangle_{1},\langle\mathcal{S}\rangle_{n-1})$, i.e.
+there is no choice ‘left vs.
+right’.
 2.2. Algebraic Spaces
-Our conventions follow that of the Stacks Project [Sta25].
+Our conventions follow that of the Stacks Project [STA26].
 Throughout we work with algebraic spaces in the ‘absolute’ setting, i.e.
-over the base scheme $\operatorname{Spec}(\mathbb{Z})$; but, of course, everything remains valid when working over an arbitrary base.
+over the base scheme $\operatorname{Spec}(\mathbb{Z})$.
+Of course, everything remains valid over an arbitrary base.
 Let $X$ be an algebraic space throughout this subsection.
-We denote by $|X|$ the set of points of $X$, which obtains a topology by giving it the quotient topology via an étale surjective morphism from a scheme; see e.g.
-[Sta25, Tag 03BT].
-The algebraic spaces that are considered in this work will always be decent in the sense of [Sta25, Tag 03I7].
+We denote by $|X|$ the set of points of $X$, endowed with the quotient topology via any étale surjective morphism from a scheme; see e.g.
+[STA26, Tag 03BT].
+The algebraic spaces that are considered in this work will always be decent in the sense of [STA26, Tag 03I7].
 This condition ensures that there is a good relationship between the topology of $|X|$ and properties of $X$ itself.
-Note that quasi-separated algebraic spaces are automatically decent; in fact, for locally Noetherian algebraic spaces being descent is equivalent to being quasi-separated [Sta25, Tag 0BB6].
+Note that quasi-separated algebraic spaces are automatically decent; in fact, for locally Noetherian algebraic spaces being descent is equivalent to being quasi-separated [STA26, Tag 0BB6].
 Defining irreducibility for algebraic spaces is a bit subtle as this notion is not étale local.
@@
 using étale surjective morphisms from schemes).
-An integral algebraic spaces, then, is one that is irreducible, reduced and decent [Sta25, Tag 0AD3].
+An algebraic spaces is integral if it’s irreducible, reduced, and decent [STA26, Tag 0AD3].
 Modules and derived categories
@@
 Consequently, it makes sense to speak of sheaves $\mathcal{O}_{X}$-modules on $X_{\acute{e}t}$; denote the (Grothendieck) Abelian category of such sheaves by $\operatorname{Mod}(X)$.
-The Abelian subcategories of quasi-coherent resp.
-coherent $\mathcal{O}_{X}$-modules are denoted by $\operatorname{Qcoh}(X)$ resp.
-$\operatorname{Coh}(X)$.
+The Abelian subcategories of quasi-coherent (resp.
+coherent) $\mathcal{O}_{X}$-modules are denoted by $\operatorname{Qcoh}(X)$ (resp.
+$\operatorname{Coh}(X)$).
 Next, let $D(X):=D(\operatorname{Mod}(X))$ denote the (unbounded) derived category of $\operatorname{Mod}(X)$ and let $D_{\operatorname{qc}}(X)$ resp.
@@
 Furthermore, let $D^{\flat}_{\#}(X):=D^{\flat}(X)\cap D_{\#}(X)$ for $\flat\in\{+,-,b,\geq n,\dots\}$ and $\in\{\operatorname{qc},\operatorname{coh}\}$, where the superscripts have to interpreted in the usual fashion.
-Any morphisms $f\colon Y\to X$ of algebraic spaces induces a morphism of ringed topoi [Sta25, Tag 03G8], which similarly as in the case of schemes leads to a derived adjunction $\mathbf{L}f^{\ast}\colon D(X)\rightleftarrows D(Y)\colon\mathbf{R}f_{\ast}$; see e.g.
-[Sta25, Tag 07A6].
+Any morphisms $f\colon Y\to X$ of algebraic spaces induces a morphism of ringed topoi [STA26, Tag 03G8], which leads to a derived adjunction $\mathbf{L}f^{\ast}\colon D(X)\rightleftarrows D(Y)\colon\mathbf{R}f_{\ast}$ as in the case of schemes; see e.g.
+[STA26, Tag 07A6].
 When $f$ is quasi-compact and quasi-separated, the derived pushforward preserves objects with quasi-coherent cohomology (see e.g.
-[Sta25, Tag 08FA]) and so the adjunction restricts to $D_{\operatorname{qc}}$, i.e.
+[STA26, Tag 08FA]) and so the adjunction restricts to $D_{\operatorname{qc}}$, i.e.
 we have an adjunction $\mathbf{L}f^{\ast}\colon D_{\operatorname{qc}}(X)\rightleftarrows D_{\operatorname{qc}}(Y)\colon\mathbf{R}f_{\ast}$.
 Perfect complexes
-Perfect complexes can be defined for any ringed site [Sta25, Tag 08G4], and so, in particular, for $X$ by looking at its small étale site.
+Perfect complexes can be defined for any ringed site [STA26, Tag 08G4], and so, in particular, for $X$ by looking at its small étale site.
 Recall that a complex is strictly perfect if it is a bounded complex with each term a direct summand of a finite free module.
@@
 The subcategory of $D_{\operatorname{qc}}(X)$ consisting of perfect complexes is denoted by $\operatorname{Perf}(X)$; this is a thick subcategory.
-When $X$ is quasi-compact and quasi-separated, the perfect complexes are exactly the compact objects of $D_{\operatorname{qc}}(X)$, see [Sta25, Tag 09M8].
+When $X$ is quasi-compact and quasi-separated, the perfect complexes are exactly the compact objects of $D_{\operatorname{qc}}(X)$, see [STA26, Tag 09M8].
 Recall that an object in a triangulated category with arbitrary direct sums is said to be compact if the covariant Hom out of that object commutes with direct sums.
-Moreover, in this case, $D_{\operatorname{qc}}(X)$ is compactly generated by a single perfect complex, see [Sta25, Tag 0AEC].
-Here $P$ compactly generating means that it is compact and any non-zero object admits a non-zero morphism from $P$, any such object is called a compact generator.
+Moreover, in this case, $D_{\operatorname{qc}}(X)$ is compactly generated by a single perfect complex, see [STA26, Tag 0AEC].
+Here, an object $P$ compactly generating means that it is compact and any non-zero object admits a non-zero morphism from some shift of $P$.
+Any such object is called a compact generator.
 It is worthwhile noting that $P$ being a compact generator for $D_{\operatorname{qc}}(X)$ is equivalent to it being a classical generator for $\operatorname{Perf}(X)$, i.e.
@@
 This follows as $D_{\operatorname{qc}}(X)$ is compactly generated, see e.g.
-[Sta25, Tag 09SR].
+[STA26, Tag 09SR].
 Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g.
-[HR17, Lemma 4.1] or [DLMRP25, Lemma 2.2].
-For the definition of the the fpqc topology, see [Sta25, Tag 03MP]; important examples of fpqc covers are fppf and so also smooth and étale covers.
+[HR17, Lemma 4.1] or [DLR+25, Lemma 2.2].
+For the definition of the the fpqc topology, see [STA26, Tag 03MP]; important examples of fpqc covers are fppf, and hence, also smooth and étale covers.
 Lemma 2.1.
@@
 Let $\overline{x}$ be a geometric point of $X$, i.e.
-$\overline{x}$ is a morphism from the spectrum of an algebraically closed field.
-For any presheaf $\mathcal{F}$ on $X_{\acute{e}t}$ one defines the stalk of $\mathcal{F}$ at $\overline{x}$ as
-$\mathcal{F}_{\bar{x}}:=\mathop{\mathrm{colim}}\nolimits_{(U,\overline{u})}\mathcal{F}(U)$
-where the colimit runs over the étale neighborhoods of $\overline{x}$ in $X$.
-Taking stalks at geometric points are exactly the points of the small étale site of $X$ (see e.g.
-[Sta25, Tag 04K6]).
+$\overline{x}$ is a morphism from the spectrum of an algebraically closed field to $X$.
+For any presheaf $\mathcal{F}$ on $X_{\acute{e}t}$, one defines the stalk of $\mathcal{F}$ at $\overline{x}$ as $\mathcal{F}_{\bar{x}}:=\mathop{\mathrm{colim}}\nolimits_{(U,\overline{u})}\mathcal{F}(U)$ where the colimit runs over the étale neighborhoods of $\overline{x}$ in $X$.
+The stalks at geometric points defined this way are exactly the stalks at points of the small étale site of $X$ (see e.g.
+[STA26, Tag 04K6]).
 The étale local ring of $X$ at $\overline{x}$ is by definition the stalk of the structure sheaf at $\overline{x}$, i.e.
@@
 This is in fact a local ring (as can be seen from the following paragraph).
-To make the latter more explicit, pick an étale neighborhood $(U,\overline{u})$ of $\overline{x}$ where $U$ a scheme. Then, from [Sta25, Tag 04KF], we have isomorphisms
+To make the latter more explicit, pick an étale neighborhood $(U,\overline{u})$ of $\overline{x}$ where $U$ a scheme. Then, from [STA26, Tag 04KF], we have isomorphisms
 (2.1) $\mathcal{O}_{X,\overline{x}}\cong\mathcal{O}_{U,\overline{u}}\cong\mathcal{O}_{U,u}^{sh},$
 where $\mathcal{O}_{U,u}^{sh}$ is the strict Henselization of $\mathcal{O}_{U,u}$ and $u\in U$ is the point at which $\overline{u}$ is centered.
-We refer to [Sta25, Tag 0BSK] for the definition of (strict) Henselizations of local rings.
-Because of Equation 2.1, for arbitrary $x\in|X|$, by abuse of notation we sometimes denote $\mathcal{O}_{X,x}^{sh}:=\mathcal{O}_{X,\overline{x}}$ where $\overline{x}$ is any geometric point lying over $x$.
+We refer to [STA26, Tag 0BSK] for the definition of (strict) Henselizations of local rings.
+Because of (2.1), for arbitrary $x\in|X|$, by abuse of notation we sometimes denote $\mathcal{O}_{X,x}^{sh}:=\mathcal{O}_{X,\overline{x}}$ where $\overline{x}$ is any geometric point lying over $x$.
 Lastly, let us briefly define the Henselian local ring at a point $x\in|X|$. Let
 $\mathcal{O}_{X,x}^{h}:=\mathop{\mathrm{colim}}\nolimits_{(U,u)}\Gamma(U,\mathcal{O}_{U})$
-where now the colimit is taken over the ‘elementary étale neighborhoods’ (see [Sta25, Tag 0BGU] for the definition).
-As in the previous paragraph one can make this more explicit in terms of a covering.
-Indeed, by [Sta25, Tag 0EMY], picking an elementary etale neighborhood $(U,u)\to(X,x)$ one has $\mathcal{O}_{X,x}^{h}\cong\mathcal{O}_{U,u}^{h}$, where $\mathcal{O}_{U,u}^{h}$ is the Henselization of the local ring $\mathcal{O}_{U,u}$ of $U$ at $u$.
+where now the colimit is taken over the ‘elementary étale neighborhoods’ (see [STA26, Tag 0BGU] for the definition).
+As in the previous paragraph, one can make this more explicit in terms of a covering.
+Indeed, by [STA26, Tag 0EMY], picking an elementary etale neighborhood $(U,u)\to(X,x)$ one has $\mathcal{O}_{X,x}^{h}\cong\mathcal{O}_{U,u}^{h}$, where $\mathcal{O}_{U,u}^{h}$ is the Henselization of the local ring $\mathcal{O}_{U,u}$ of $U$ at $u$.
 In particular,
 (2.2) $(\mathcal{O}_{X,x}^{h})^{sh}=\mathcal{O}_{X,x}^{sh},$
-see e.g. [Sta25, Tag 0EMZ].
+see e.g. [STA26, Tag 0EMZ].
 Residue fields
 Pick a point $x\in|X|$.
-When $X$ is decent there exists, by [Sta25, Tag 03K4], a unique monomorphism $\operatorname{Spec}(k)\to X$ representing $x$.
+When $X$ is decent there exists, by [STA26, Tag 03K4], a unique monomorphism $\operatorname{Spec}(k)\to X$ representing $x$.
 The field $k$ appearing is called the residue field of $X$ at $x$ and is denoted $\kappa(x)$.
@@
 Recall that $X$ is regular if there exists an étale surjective morphism from a regular scheme, see e.g.
-[Sta25, Tag 03E5].
-Furthermore, a point $x\in|X|$ is regular if there exists an étale morphism $f\colon U\to X$ from a scheme and a point $u\in U$ such that $f(u)=x$ and the stalk $\mathcal{O}_{U,u}$ is a regular local ring; this is well-defined by [Sta25, Tag 0AH7].
+[STA26, Tag 03E5].
+Furthermore, a point $x\in|X|$ is regular if there exists an étale morphism $f\colon U\to X$ from a scheme and a point $u\in U$ such that $f(u)=x$ and the stalk $\mathcal{O}_{U,u}$ is a regular local ring; this is well-defined by [STA26, Tag 0AH7].
 It is straightforward to see that an algebraic space is regular if and only if all of its points are regular.
-Moreover, by [Sta25, Tag 0AHA], a point $x\in|X|$ is regular if and only if the étale local ring $\mathcal{O}_{X,\overline{x}}$ is a regular local ring for any (equivalently some) geometric point $\overline{x}$ lying over $x$.
-When $X$ is a Noetherian algebraic space, it follows from [DLMRP25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$.
-Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLMRP25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e.
+Moreover, by [STA26, Tag 0AHA], a point $x\in|X|$ is regular if and only if the étale local ring $\mathcal{O}_{X,\overline{x}}$ is a regular local ring for any (equivalently some) geometric point $\overline{x}$ lying over $x$.
+When $X$ is a Noetherian algebraic space, it follows from [DLR+25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$.
+Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLR+25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e.
 $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.
@@
 We say a morphism of integral algebraic spaces $f\colon Y\to X$ is
-  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [Sta25, Tag 0ACU] for some characterizations and
-  • an alteration if it is proper and generically finite, see e.g. [Sta25, Tag 0AD7] for some characterizations.
+  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [STA26, Tag 0ACU] for some characterizations and
+  • an alteration if it is proper and generically finite, see e.g. [STA26, Tag 0AD7] for some characterizations.
 Usually, in this work the target (and hence, source) of the morphism will be Noetherian.
@@
 The modifications we will be interested in below arise as blowups.
-See [Sta25, Tag 085P] some background on blowups of algebraic spaces.
+See [STA26, Tag 085P] for some background on blowups of algebraic spaces.
 Let $f\colon Y\to X$ now be more generally any morphism of quasi-compact quasi-separated algebraic spaces. Extending [LN07] to algebraic spaces we obtain the following.
@@
 These notions coincide, for example, for a proper morphisms of Noetherian schemes [LN07, Theorem 1.2], but differ in general.
-3. Results
-To start, we note an ‘obvious’ lemma which is recorded for convenience.
-Lemma 3.1.
-Let $f\colon Y\to X$ be a morphisms of quasi-compact quasi-separated algebraic spaces with $Y$ nonempty.
-Then $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is nonzero; in particular, $\mathbf{R}f_{\ast}G$ is nonzero for any compact generator $G$ of $D_{\operatorname{qc}}(Y)$.
-Proof.
-The second claim follows from the first as $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is finitely built from $\mathbf{R}f_{\ast}G$ (i.e.
-$\mathcal{O}_{Y}\in\langle G\rangle$ and so also after applying $\mathbf{R}f_{\ast}$).
-To see the first claim, observe that it follows from pullback/pushforward adjunction that
-$0$ $\not=H^{0}(Y,\mathcal{O}_{Y})$
-$\cong\operatorname{Ext}^{0}(\mathcal{O}_{Y},\mathcal{O}_{Y})$
-$\cong\operatorname{Ext}^{0}(\mathbf{L}f^{\ast}\mathcal{O}_{X},\mathcal{O}_{Y})$
-$\cong\operatorname{Ext}^{0}(\mathcal{O}_{X},\mathbf{R}f_{\ast}\mathcal{O}_{Y})$
-and so $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is nonzero. ∎
-The next should be compared to [Sta25, Tag 00OC] and gives a ‘spacy’ version of this.
+3. Quasi-perfect blowups detect regularity
+This section shows a relationship between regularity and having ‘enough’ quasi-perfect ‘probing morphisms’. The following should be compared to [STA26, Tag 00OC] and gives a ‘spacy’ version of this.
 Lemma 3.1.
@@
 a Noetherian) algebraic space.
-Suppose $x\in|X|$ is a closed point such that $i_{x,*}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}$ is perfect, where $i_{x}\colon{\operatorname{Spec}}(\kappa(x))\to X$ denotes the residue field of $X$ at $x$.
+Suppose $x\in|X|$ is a closed point such that $i_{x,\ast}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}$ is perfect, where $i_{x}\colon{\operatorname{Spec}}(\kappa(x))\to X$ denotes the residue field of $X$ at $x$.
 Then $x$ is a regular point of $X$.
 Proof.
-Let $(U,u)\to(X,x)$ be the ‘elementary étale neighborhood’ of [Sta25, Tag 0BBP], i.e.
+Let $(U,u)\to(X,x)$ be the ‘elementary étale neighborhood’ of [STA26, Tag 0BBP], i.e.
 $U$ is an affine scheme, $u$ is the only point of $U$ lying over $x$, and the induced homomorphism of residue fields $\kappa(x)\to\kappa(u)$ is an isomorphism.
@@
 It follows from categorical nonsense that $U_{x}\to U$ is a monomorphism (as $\operatorname{Spec}(\kappa(x))\to X$ is).
-Moreover, as $U_{x}\to\operatorname{Spec}\kappa(x)$ is étale, it is a disjoint union of fields.
+Moreover, as $U_{x}\to\operatorname{Spec}(\kappa(x))$ is étale, it is a disjoint union of fields.
 However, as each field corresponds to a distinct point lying over $x$ (they all go to different points as monomorphisms of schemes are injective), $U_{x}$ is the spectrum of a field.
@@
 In particular, it also follows that $u$ is necessarily a closed point as $i_{x}$ is a closed immersion by e.g.
-[Sta25, Tag 0H1R & Tag 0H1U].
-Next, using e.g. flat base change, one sees
-$f^{*}i_{x,*}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}=i_{u,*}\mathcal{O}_{\operatorname{Spec(\kappa(u))}}$
+[STA26, Tag 0H1R & Tag 0H1U].
+Next, using flat base change, one sees
+$f^{\ast}i_{x,\ast}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}=i_{u,\ast}\mathcal{O}_{\operatorname{Spec(\kappa(u))}}$
 and so also perfect.
 Taking the stalk at $u$, one has the residue field $\kappa(u)$ is a perfect $\mathcal{O}_{U,u}$-module.
-Thus, $u$ is a regular point by [Sta25, Tag 00OC] and so $x$ is a regular point of $X$ by definition.
+Thus, $u$ is a regular point by [STA26, Tag 00OC] and so $x$ is a regular point of $X$ by definition.
 ∎
-Theorem 3.3.
-A Noetherian algebraic space $X$ is regular if, and only if, the blowup of $X$ along any closed point is a quasi-perfect morphism.
+Lemma 3.2.
+Let $X$ be Noetherian algebraic space $X$. If the blowup of $X$ along a closed point $p\in|X|$ induces a quasi-perfect morphisms, then $p$ is regular.
 Proof.
-If $X$ is regular, then for any proper $f$ we have
-$\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\subseteq D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X).$
-Hence, any blowup is quasi-perfect as these are proper.
-For the converse, it suffices to show that any closed point is regular (see e.g.
-[DLMRP25, Lemma 3.2]).
-So, let $p\in|X|$ be a closed point and denote by $i\colon\operatorname{Spec}(\kappa(p))\to X$ its associated residue field.
-As the point is closed, $i$ is a closed immersion; see e.g.
-[Sta25, Tag 0H1R & Tag 0H1U].
+Denote by $i\colon\operatorname{Spec}(\kappa(p))\to X$ the residue field associated to our closed point $p$; this is a closed immersion, see e.g.
+[STA26, Tag 0H1R & Tag 0H1U].
 Consider the fibered square
 […]
-where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor. The hypothesis tells us that $f$ is a quasi-perfect morphism.
-As the exceptional divisor $E$ is an effective Cartier divisor on $X^{\prime}$, we have a short exact sequence
-$0\to\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to 0$
-from which we obtain a distinguished triangle in $D^{b}_{\operatorname{coh}}(X^{\prime})$,
-(3.2) $\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to\mathcal{O}_{X^{\prime}}(-E)[1].$
-Furthermore, as $\mathcal{O}_{X^{\prime}}(-E)$ is an invertible sheaf, and so perfect, we have $j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X^{\prime})$.
-Hence, as $f$ is quasi-perfect by assumption, we know that $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$.
-Moreover, from the commutativity of Equation 3.1 and Lemma 3.1, we see that
+where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor.
+By definition of blowing up the exceptional divisor $E$ is an effective Cartier divisor, and so it follows from the distinguished triangle
+$\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to\mathcal{O}_{X^{\prime}}(-E)[1].$
+that $j_{\ast}\mathcal{O}_{E}$ is perfect.
+Hence, as $f$ is quasi-perfect by assumption, $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$ is perfect too.
+To conclude, observe from the commutativity of Diagram˜3.1 that
 $0$ $\not=\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$
@@
 $\cong\text{sum of shifts of }i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$
-It follows that $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect.
-Hence $p$ is regular by Lemma 3.2 and we are done.
+Hence $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect.
+Thus, $p$ is regular by Lemma˜3.1, and we are done.
 ∎
-Remark 3.4.
-It is worthwhile to stress the following: the blowup at a closed point being quasi-perfect implies that that point is regular. Indeed, this is essentially the content of the above proof.
+Proposition 3.3.
+A Noetherian algebraic space $X$ is regular if and only if the blowup of $X$ along any closed point is a quasi-perfect morphism.
+Proof.
+If $X$ is regular, then for any proper $f$ we have
+$\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\subseteq D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X).$
+Hence, any blowup is quasi-perfect as these are proper. The converse follows from Lemma˜3.2 as it suffices to check regularity at closed points (see e.g. [DLR+25, Lemma 3.2]). ∎
 Lemma 3.4.
@@
 The blowup is reduced, see e.g.
-[Sta25, Tag 085W], as the latter is an étale local property.
+[STA26, Tag 085W], as the latter is an étale local property.
 Moreover, since blowups are representable, it follows that $X^{\prime}$ is decent (see e.g.
-[Sta25, Tag 0ABT]).
+[STA26, Tag 0ABT]).
 Thus, it remains to show that $|X^{\prime}|$ is irreducible, which is more subtle as this notion is not étale local.
@@
 We have that $f^{-1}(Z)$ is an effective Cartier divisor on $X^{\prime}$ and that $f$ induces an isomorphism $f^{-1}(X\setminus Z)=X\setminus Z$; see e.g.
-[Sta25, Tag 085T].
+[STA26, Tag 085T].
 When $|X|$ is irreducible so is the open subset $|X\setminus Z|$ and so $|f^{-1}(X\setminus Z)|=|X^{\prime}\setminus f^{-1}(Z)|$ is also.
-However, $f^{-1}(Z)$ being effective Cartier implies that the open subspace $X^{\prime}\setminus f^{-1}(Z)$ is scheme-theoretically dense in $X^{\prime}$, see [Sta25, Tag 083S].
+However, $f^{-1}(Z)$ being effective Cartier implies that the open subspace $X^{\prime}\setminus f^{-1}(Z)$ is scheme-theoretically dense in $X^{\prime}$, see [STA26, Tag 083S].
 In particular, the closure of $|X^{\prime}\setminus f^{-1}(Z)|$ (in $|X^{\prime}|$) is $|X^{\prime}|$ and so the latter is also irreducible.
-To see that $f$ is a modification when the ideal is of finite type, note that in this case $f$ is proper by [Sta25, Tag 085Z]. ∎
+To see that $f$ is a modification when the ideal is of finite type, note that in this case $f$ is proper by [STA26, Tag 085Z]. ∎
 Corollary 3.5.
-An integral Noetherian algebraic space $X$ is regular if, and only if, every alteration (resp. modification) of $X$ is quasi-perfect.
+An integral Noetherian algebraic space $X$ is regular if and only if every alteration (resp. modification) of $X$ is quasi-perfect.
 Proof.
-This follows immediately from Theorem 3.3.
-Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma 3.5.
+This follows immediately from Lemma˜3.2.
+Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma˜3.4.
 Moreover, as modifications are alterations the case of alterations also follows.
@@
 In the special case where there exists an alteration or modification $f\colon Y\to X$ of algebraic spaces where $Y$ is regular and $f$ quasi-perfect, then one can show $X$ regular more easily using ‘proper descent’.
-For example, this follows from [HLLP25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity.
-So, the novelty of Corollary 3.6 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
-Next come some lemmas that allow one to check quasi-perfectness using flat covers.
+For example, this follows from [HLL+25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity.
+So, the novelty of Corollary˜3.5 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
+4. On proper quasi-perfect morphisms
+Here we study quasi-perfect morphisms which are proper. We start with a few lemmas that allow one to check quasi-perfectness using flat covers.
 Lemma 4.1.
-A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if, and only if, $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
+A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if and only if $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
 Proof.
@@
   1. (1)
-    If $f$ is quasi-perfect and $s$ is affine, then $f^{\prime}$ is quasi-perfect.
+    If $f$ is quasi-perfect and $s$ is quasi-affine, then $f^{\prime}$ is quasi-perfect.
   2. (2)
@@
 Proof.
-To show the first claim, first note that when $s$ is an affine morphism so is its base change $s^{\prime}$.
-Consequently, $\mathbf{R}s^{\prime}_{\ast}=s^{\prime}_{\ast}$ is conservative (this statement is local on the target and so reduces to the case of schemes) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g.
-[Sta25, Tag 0E4R]).
+To show the first claim, first note that when $s$ is an quasi-affine morphism so is its base change $s^{\prime}$.
+Consequently, $\mathbf{R}s^{\prime}_{\ast}$ is conservative (see e.g.
+[HR17, Corollary 2.8]) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g.
+[STA26, Tag 0E4R]).
 Thus, as $f$ is assumed quasi-perfect, $\mathbf{R}f_{\ast}Q$ is perfect and so also is its pullback $s^{\ast}\mathbf{R}f_{\ast}Q$.
@@
 (4.2) $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q\cong s^{\ast}\mathbf{R}f_{\ast}Q.$
-Hence, Lemma 3.8 implies $f^{\prime}$ is a quasi-perfect morphism.
+Hence, Lemma˜4.1 implies $f^{\prime}$ is a quasi-perfect morphism.
 For the second claim, assume now that $f$ is surjective and $f^{\prime}$ is quasi-perfect.
-Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation 3.4.
+Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation˜4.2.
 As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact.
-Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma 2.1.
-By Lemma 3.8, we have that $f$ is quasi-perfect, which completes the proof.
+Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma˜2.1.
+By Lemma˜4.1, we have that $f$ is quasi-perfect, which completes the proof.
 ∎
@@
 Proof.
-The forward direction follows from Lemma 3.9.
-For the converse, denote the base change of $f$ along $s$ as in Equation 3.3.
+The forward direction follows from Lemma˜4.2.
+For the converse, denote the base change of $f$ along $s$ as in Diagram˜4.1.
 Again, choose a compact generator $Q$ of $Y$.
-By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation 3.4, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect.
-However, as $s$ corresponds to a faithfully flat ring map, it follows from [Sta25, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect.
-Thus, Lemma 3.8 shows that $f$ is quasi-perfect.
+By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation˜4.2, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect.
+However, as $s$ corresponds to a faithfully flat ring map, it follows from [STA26, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect.
+Thus, Lemma˜4.1 shows that $f$ is quasi-perfect.
 ∎
-Using Lemma 3.9, we obtain an alternative proof of the fact that regularity can be checked fpqc locally.
+Using Lemma˜4.2, we obtain an alternative proof of the fact that regularity can be checked fpqc locally. Before doing so, we recall a fact whose proof we add for lack of finding a reference.
+Lemma 4.4.
+Let $X$ be a quasi-separated algebraic space. Then any morphism from an affine scheme to $X$ is quasi-affine.
+Proof.
+Suppose $U\to X$ is a map from an affine scheme.
+So show it is affine, by definition, we need to show, for any affine scheme $V$ and morphisms $V\to X$, that $U\times_{X}V$ is quasi-affine.
+As the latter is just the pullback of the morphisms $U\times V\to X\times X$ along the diagonal $X\to X\times X$, this follows from the fact the diagonal of $X$ is quasi-affine, see e.g.
+[STA26, Tag 03HK & Tag 02LR].
+∎
 Corollary 4.5.
-Let $X$ be a separated Noetherian algebraic space. If there is an fpqc cover $\{f_{i}\colon U_{i}\to X\}_{i\in I}$ where each $U_{i}$ is regular, then $X$ is regular.
+Let $X$ be a Noetherian algebraic space. If there is an fpqc cover $\{f_{i}\colon U_{i}\to X\}_{i\in I}$ where each $U_{i}$ is regular, then $X$ is regular.
 Proof.
@@
 take étale surjective morphisms from affine schemes).
-Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is affine by [Sta25, Tag 08GB] as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme.
-To show $X$ is regular, it suffices by Theorem 3.3 to show that blowups along nonzero ideal sheaves are quasi-perfect.
+Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is quasi-affine by Lemma˜4.4 as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme.
+To show $X$ is regular, it suffices by Lemma˜3.2 to show that blowups along nonzero ideal sheaves are quasi-perfect.
 However, as blowups commute with flat base change, see e.g.
-[Sta25, Tag 0805], we have that the base change to $U$ is again such a blow-up.
+[STA26, Tag 0805], we have that the base change to $U$ is again such a blow-up.
 Consequently, as $U$ is regular, it is quasi-perfect.
-Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma 3.9, showing the claim.
+Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma˜4.2, showing the claim.
 ∎
@@
 Indeed, it follows from the existence of an fppf cover as in the statement that $X$ is locally Noetherian (as this is fppf local and regular algebraic spaces are always locally Noetherian) and hence Noetherian.
-There are three important cases of faithfully flat morphisms that we will apply Lemma 3.10 to.
+There are three important cases of faithfully flat morphisms that we will apply Lemma˜4.3.
 Example 4.7.
@@
 There are local ring morphisms $R\to\widehat{R}$ and $R\to R^{h}\to R^{sh}$ which are flat and consequently faithfully flat; see e.g.
-[Sta25, Tag 00MC and Tag 07QM].
+[STA26, Tag 00MC and Tag 07QM].
 (As an aside, there is also a flat local ring morphism $R^{h}\to\widehat{R}$, but we do not need it).
@@
 Theorem 4.8.
-Let $f\colon Y\to X$ be a proper morphism of algebraic spaces with $X$ separated and Noetherian.
-Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$.
-Then the following are equivalent:
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
   1. (1)
@@
 ($\star$) If the base change of $f$ to $\mathcal{O}_{U,p}$ is quasi-perfect for all $p\in U$, then $f$ is quasi-perfect.
-The proof of ($\star$3) is similar to that of Lemma 3.9; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work.
+The proof of ($\star$4) is similar to that of Lemma˜4.2; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work.
 Pick a point $p\in U$ and look at the fibered square
@@
 As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper.
-Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma 2.1 and $f$ is quasi-perfect by Lemma 3.8.
+Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma˜2.1 and $f$ is quasi-perfect by Lemma˜4.1.
 ∎
 Corollary 4.9.
-Let $f\colon Y\to X$ be a proper morphism of algebraic spaces with $X$ separated and Noetherian and let $A_{\#}$ and $s\colon U\to X$ be as in Theorem 3.14. Then the locus
-(3.5) $S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
+$S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
 is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
 Proof.
-By the proof of Theorem 3.14 (i.e. using Lemma 3.10) we know that
-$S_{X}:=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
+By the proof of Theorem˜4.8 (i.e. using Lemma˜4.3) we know that
+$S_{X}$= $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
 $=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{h}\textrm{ is quasi-perfect}\right\}$
 and
-$S_{U}:=$ $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
+$S_{U}$= $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
 $=$ $\left\{u\in U\mid\text{base change to }\widehat{\mathcal{O}}_{U,u}\textrm{ is quasi-perfect}\right\}.$
@@
 (4.3) $s^{-1}(S_{X})=S_{U}.$
-To this end, observe by Equation 2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$.
-As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma 3.10.
+To this end, observe by Equation˜2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$.
+As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma˜4.3.
 Therefore,
@@
 $=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}=S_{U}.$
-As a consequence of Equation 3.6 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open.
+As a consequence of Equation˜4.3 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open.
 Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
@@
 Remark 4.10.
-One can think of the subset $S$ appearing in Corollary 3.15 as the ‘quasi-perfect locus of $f$’.
+One can think of the subset $S$ appearing in Corollary˜4.9 as the ‘quasi-perfect locus of $f$’.
 It follows from the first part of the proof that in general, i.e.
-regardless of properness of $f$, one can compute it étale locally (Equation 3.6) and that the specific choice of $A_{\#}$ does not matter.
+regardless of properness of $f$, one can compute it étale locally (Equation˜4.3) and that the specific choice of $A_{\#}$ does not matter.
 Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
@@
 Proposition 4.11.
-Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [Sta25, Tag 04NA] for the relevant definition.
+Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [STA26, Tag 04NA] for the relevant definition.
 whose étale local ring is reduced (e.g.
@@
   • As $X$ admits an open dense subscheme, which in particular contains $\eta$, see e.g.
-[Sta25, Tag 06NH], we may assume that $X$ is a scheme.
+[STA26, Tag 06NH], we may assume that $X$ is a scheme.
 Moreover, by choosing any affine open neighborhood $U$ of $\eta$ and by restricting to $f^{-1}(U)\to U$, we may assume $X$ is affine and, in particular, separated.
-  • Next, let $S$ denote the open subset from Corollary 3.15; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty.
-It follows from Theorem 3.14 that the restriction $f^{-1}(S)\to S$ is quasi-perfect.
+  • Next, let $S$ denote the open subset from Corollary˜4.9; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty.
+It follows from Theorem˜4.8 that the restriction $f^{-1}(S)\to S$ is quasi-perfect.
 So we may assume $f$ is quasi-perfect.
@@
 In order to finish the proof, is suffices to show that $X$ is regular.
-This follows from ‘proper descent’, as already indicated in Remark 3.7; but we give a self-contained proof, more in line with that of Theorem 3.3, for the convenience of the reader.
+This follows from ‘proper descent’, as already indicated in Remark˜3.6; but we give a self-contained proof, more in line with that of Lemma˜3.2, for the convenience of the reader.
 So, let $p\in X$ be a closed point and denote the associated closed immersion by $i\colon\mathrm{Spec}(\kappa(p))\to X$. Consider the fibered square

### v3

@@
 Abstract.
-This note records two results concerning quasi-perfect morphisms between Noetherian algebraic spaces.
-First, we give a new characterization of regular Noetherian algebraic spaces as those for which blowups at closed points are quasi-perfect.
-Secondly, we study the local behavior of quasi-perfect proper morphisms.
-In particular, we show that quasi-perfectness of a proper morphism can be detected at their étale local rings, completions, and (strict) Henselizations.
-As a corollary, the locus of points where a proper morphism is quasi-perfect is Zariski open.
+This note is concerned with quasi-perfect morphisms between Noetherian algebraic spaces.
+In particular, we study the local behavior of quasi-perfect proper morphisms.
+We show that quasi-perfectness of a proper morphism can be detected at the étale local rings of points of the target, as well as their completions and (strict) Henselizations.
+As a corollary, we obtain that the locus of points where a proper morphism is quasi-perfect is Zariski open.
 Key words and phrases:
-Derived category, quasi-perfect morphism, blowup, regularity, generation (for triangulated categories)
+Algebraic spaces, quasi-perfect morphism, generation (for triangulated categories), regularity
 2020 Mathematics Subject Classification:
-14A30 (primary), 14B05, 18G80, 13D09
+14A30 (primary), 13D09, 14B05, 18G80
 1. Introduction
-Given a quasi-compact quasi-separated morphism $f\colon Y\to X$ of schemes, Neeman’s Brown Representability Theorem provides a right adjoint $f^{\times}$ to the derived pushforward $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$.
+Given a quasi-compact and quasi-separated morphism $f\colon Y\to X$ of schemes, Neeman’s Brown Representability Theorem provides a right adjoint $f^{\times}$ to the derived pushforward $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$, the derived category of complexes with quasi-coherent cohomology.
 While $f^{\times}$ preserves small coproducts for complexes that are uniformly homologically bounded below, this property fails in general.
-Motivated by [VER69, Lemma 1 & Corollary 2], Lipman–Neeman [LN07] introduced the notion of quasi-perfect morphisms for quasi-compact quasi-separated schemes to remedy this.
-The defining property is that $f^{\times}$ preserves small coproducts on all of $D_{\operatorname{qc}}$ or, equivalently, that $\mathbf{R}f_{\ast}$ preserve perfect complexes; see [LN07, Theorem 1.2 & Proposition 2.1].
-Whilst there are numerous examples of quasi-perfect morphisms—for instance, any proper morphisms of finite Tor dimension between Noetherian schemes—quasi-perfectness generally fails.
-For example, this is the case for closed immersions into a non-regular Noetherian scheme.
+Motivated by [VER69, Lemma 1 & Corollary 2], Lipman–Neeman [LN07] introduced the notion of a quasi-perfect morphism to remedy this.
+The defining property is that $f^{\times}$ preserves small coproducts on all of $D_{\operatorname{qc}}$ which is equivalent to $\mathbf{R}f_{\ast}$ preserving perfect complexes; see [LN07, Theorem 1.2 & Proposition 2.1].
+There are numerous examples of quasi-perfect morphisms, e.g.
+any proper morphisms of finite Tor dimension between Noetherian schemes.
+However, quasi-perfectness generally fails; for example, for closed immersions into nonregular Noetherian schemes.
 More recent developments have recast these ideas in a categorical framework.
@@
 In particular, the local behavior of quasi-perfectness remains less understood; for instance, how quasi-perfectness at the local rings of the target influences quasi-perfectness globally.
-In this short note we make two novel observations regarding quasi-perfect morphisms between algebraic spaces. The first observation is how one can use these morphisms to characterize regularity.
-Proposition (Proposition˜3.3).
-A Noetherian algebraic space $X$ is regular if and only if the blowup of $X$ along any closed point is a quasi-perfect morphism.
-To the best of our knowledge, this characterization has not appeared before in the literature even for schemes.
-The simple, but crucial, observation is that a closed point on a Noetherian algebraic space is regular if and only if the blowup at that point is quasi-perfect.
-This follows from the fact that perfectness of the skyscraper sheaf at a closed point detects regularity of that point; we prove this in Lemma˜3.1, which seems to not have been written down for algebraic spaces before and might be of independent interest.
-Furthermore, as an application of Proposition˜3.3, we give a new proof that regularity descends along fpqc covers (see Corollary˜4.5).
-The second observation concerns the local behavior of quasi-perfectness for proper morphisms. This is summarized in the following result.
-Theorem (Theorem˜4.8).
+In this note, we study the local behavior of quasi-perfectness for proper morphisms between algebraic spaces.
+Theorem (Theorem˜3.4).
 Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
@@
 Furthermore, if $s\colon U\to X$ is an étale surjective morphism from a quasi-compact scheme, one can take $A_{\#}\in\{\mathcal{O}_{U,\#},\widehat{\mathcal{O}}_{U,\#}\}$ and let $p$ run over all points in $U$.
-Theorem˜4.8 provides a mechanism for detecting quasi-perfectness by reducing the question to targets that are local rings.
-The proof relies on several auxiliary lemmas, potentially of independent interest, which allow quasi-perfectness to be checked using specific classes of flat covers (see Lemmas˜4.2 and 4.3).
-As a corollary, we show that the locus of points $p$, for which the base change to one of the classes of local rings $A_{p}$ from Theorem˜4.8 is quasi-perfect, forms a Zariski open subset.
-Corollary (Corollary˜4.9).
-Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
+To the best of our knowledge, the observed behavior is new even for schemes.
+Theorem˜3.4 provides a mechanism for detecting quasi-perfectness by reducing the question to targets that are local rings.
+The proof relies on several auxiliary lemmas, potentially of independent interest, that allow for quasi-perfectness to be checked using specific classes of flat covers (see Lemmas˜3.2 and 3.2).
+As a corollary, we have the following.
+Corollary (Corollary˜3.6).
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜3.4. Then the locus
 $S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
 is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
-One can think of $S$ as the ‘the quasi-perfect locus of $f$’.
+One can think of the open subset $S$ as the ‘the quasi-perfect locus of $f$.
 The fact it is always open for proper morphisms appears to be new even for schemes.
-As an application, we show that the property that the regular locus contains a nonempty open subset (i.e.
-that $X$ is JJ-0) descends along proper surjective morphisms under mild hypotheses, which are satisfied if, for example, the target is reduced (see Proposition˜4.11).
-Finally, we comment on the restriction to proper morphisms in Theorem˜4.8.
-In view of Proposition˜3.3, this is mild since the morphisms appearing there are proper.
-In fact, for a separated morphism of finite type between Noetherian schemes, quasi-perfectness already implies properness (see [NEE21, Proposition 0.20]); so this is quite a mild restriction.
+As an application, we show the following.
+Proposition.
+Let $X$ be a reduced Noetherian algebraic space.
+Suppose further that $f\colon Y\to X$ is a proper surjective morphism.
+Then the regular locus of $X$ contains a nonempty open subset (i.e.
+is $J\textrm{-}0$) whenever $Y$ has this property.
+Again, to the best of our knowledge, this is new even in the case of schemes.
+In fact, we prove a slightly more general statement, see Proposition˜3.8.
+Furthermore, in Appendix˜A, we record a characterization for regularity of Noetherian algebraic spaces in terms of quasi-perfectness of blowups.
+Finally, we comment on the restriction to proper morphisms in Theorem˜3.4.
+In practice, it is a mild restriction.
+Indeed, for a separated morphism of finite type between Noetherian schemes, quasi-perfectness already implies properness (see [NEE21, Proposition 0.20]).
 The properness assumption is required in our arguments to employ local-to-global techniques involving bounded coherent objects (e.g.
 [AIL10, Theorem 4.1]).
-We expect the more general results to be useful in future work on singularities of algebraic spaces, a topic that has recently been gaining traction.
-For example, recently the minimal model program has been established for excellent algebraic spaces in characteristic zero [LM22].
+We expect the more general results to be useful in future work on singularities of algebraic spaces.
+This a topic that has recently been gaining traction.
+For example, recently the minimal model program has been established for excellent algebraic $\mathbb{Q}$-spaces [LM22].
 Acknowledgments.
@@
 Lank was supported under the ERC Advanced Grant 101095900-TriCatApp.
-The authors thank Leovigildo Alonso Tarrío, Fei Peng, and Karl Schwede for looking at an earlier version of this work.
+The authors thank Leovigildo Alonso Tarrío, Bhargav Bhatt, Linquan Ma, Fei Peng, and Karl Schwede for discussions and/or comments.
 Lank would like to dedicate this work to Boz, whose joyous and warm memories left behind will be forever remembered.
@@
 An algebraic spaces is integral if it’s irreducible, reduced, and decent [STA26, Tag 0AD3].
-Modules and derived categories
+Derived categories and functors
 Let $X_{\acute{e}t}$ denote the small étale site of $X$.
@@
 Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g.
-[HR17, Lemma 4.1] or [DLR+25, Lemma 2.2].
+[HR17, Lemma 4.1] or [DLM+25, Lemma 2.2].
 For the definition of the the fpqc topology, see [STA26, Tag 03MP]; important examples of fpqc covers are fppf, and hence, also smooth and étale covers.
@@
 Moreover, by [STA26, Tag 0AHA], a point $x\in|X|$ is regular if and only if the étale local ring $\mathcal{O}_{X,\overline{x}}$ is a regular local ring for any (equivalently some) geometric point $\overline{x}$ lying over $x$.
-When $X$ is a Noetherian algebraic space, it follows from [DLR+25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$.
-Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLR+25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e.
+When $X$ is a Noetherian algebraic space, it follows from [DLM+25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$.
+Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLM+25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e.
 $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.
 2.3. Morphisms
-We say a morphism of integral algebraic spaces $f\colon Y\to X$ is
-  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [STA26, Tag 0ACU] for some characterizations and
-  • an alteration if it is proper and generically finite, see e.g. [STA26, Tag 0AD7] for some characterizations.
-Usually, in this work the target (and hence, source) of the morphism will be Noetherian.
-Note that, for either type, the morphism above is, in fact, surjective.
-This follows as they are by definition dominant (this is implicit in generically finite/an isomorphism) and a dominant proper morphism is surjective.
-The modifications we will be interested in below arise as blowups.
-See [STA26, Tag 085P] for some background on blowups of algebraic spaces.
 Let $f\colon Y\to X$ now be more generally any morphism of quasi-compact quasi-separated algebraic spaces. Extending [LN07] to algebraic spaces we obtain the following.
@@
 These notions coincide, for example, for a proper morphisms of Noetherian schemes [LN07, Theorem 1.2], but differ in general.
-3. Quasi-perfect blowups detect regularity
-This section shows a relationship between regularity and having ‘enough’ quasi-perfect ‘probing morphisms’. The following should be compared to [STA26, Tag 00OC] and gives a ‘spacy’ version of this.
+3. Proofs
+We prove our main results. To start, we need a few lemmas that allow one to check quasi-perfectness using flat covers.
 Lemma 3.1.
+A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if and only if $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
+Proof.
+If $f$ is quasi-perfect, it is clear that $\mathbf{R}f_{\ast}P\in\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$.
+For, the converse direction, let $P$ be compact generator on $Y$.
+Our hypothesis is $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$.
+The desired claim then follows from the following string of inclusions:
+$\mathbf{R}f_{\ast}\operatorname{Perf}(Y)=\mathbf{R}f_{\ast}\langle P\rangle\subseteq\langle\mathbf{R}f_{\ast}P\rangle\subseteq\operatorname{Perf}(X).\qed$
+Lemma 3.2.
+Let $f\colon Y\to X$ be a morphism of quasi-compact quasi-separated algebraic spaces and let $s\colon U\to X$ be a flat morphism (from an algebraic space). Consider the fibered square
+[…]
+  1. (1)
+    If $f$ is quasi-perfect and $s$ is quasi-affine, then $f^{\prime}$ is quasi-perfect.
+  2. (2)
+    If $s$ is surjective and $f^{\prime}$ is quasi-perfect, then $f$ is quasi-perfect.
+Proof.
+To show the first claim, first note that when $s$ is an quasi-affine morphism so is its base change $s^{\prime}$.
+Consequently, $\mathbf{R}s^{\prime}_{\ast}$ is conservative (see e.g.
+[HR17, Corollary 2.8]) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g.
+[STA26, Tag 0E4R]).
+Thus, as $f$ is assumed quasi-perfect, $\mathbf{R}f_{\ast}Q$ is perfect and so also is its pullback $s^{\ast}\mathbf{R}f_{\ast}Q$.
+From flat base change, it follows that
+(3.2) $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q\cong s^{\ast}\mathbf{R}f_{\ast}Q.$
+Hence, Lemma˜3.1 implies $f^{\prime}$ is a quasi-perfect morphism.
+For the second claim, assume now that $s$ is surjective and $f^{\prime}$ is quasi-perfect.
+Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation˜3.2.
+As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact.
+Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma˜2.1.
+By Lemma˜3.1, we have that $f$ is quasi-perfect, which completes the proof.
+∎
+There are three important cases of faithfully flat morphisms that we will apply Lemma˜3.2.
+Example 3.3.
+Let $(R,\mathfrak{m})$ be a Noetherian local ring. To this one can associate a few other Noetherian local rings:
+  1. (1)
+    the $\mathfrak{m}$-adic completion of $R$, denoted $\widehat{R}$,
+  2. (2)
+    the Henselization of $R$, denoted $R^{h}$, and
+  3. (3)
+    the strict Henselization of $R$, denoted $R^{sh}$.
+There are local ring morphisms $R\to\widehat{R}$ and $R\to R^{h}\to R^{sh}$ which are flat and consequently faithfully flat; see e.g.
+[STA26, Tag 00MC and Tag 07QM].
+(As an aside, there is also a flat local ring morphism $R^{h}\to\widehat{R}$, but we do not need it).
+We now move to our main result.
+Theorem 3.4.
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
+  1. (1)
+    $f$ is quasi-perfect,
+  2. (2)
+    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in|X|$.
+Furthermore, if $s\colon U\to X$ is an étale surjective morphism from a quasi-compact scheme, one can take $A_{\#}\in\{\mathcal{O}_{U,\#},\widehat{\mathcal{O}}_{U,\#}\}$ and let $p$ run over all points in $U$.
+Before proving this, we recall a fact whose proof we add for lack of finding a reference.
+Lemma 3.5.
+Let $X$ be a quasi-separated algebraic space. Then any morphism from an affine scheme to $X$ is quasi-affine.
+Proof.
+Suppose $U\to X$ is a map from an affine scheme.
+So show it is affine, by definition, we need to show, for any affine scheme $V$ and morphisms $V\to X$, that $U\times_{X}V$ is quasi-affine.
+As the latter is just the pullback of the morphisms $U\times V\to X\times X$ along the diagonal $X\to X\times X$, this follows from the fact the diagonal of $X$ is quasi-affine, see e.g.
+[STA26, Tag 03HK & Tag 02LR].
+∎
+Proof of Theorem˜3.4.
+[…]
+[…]
+($\star$) If the base change of $f$ to $\mathcal{O}_{U,p}$ is quasi-perfect for all $p\in U$, then $f$ is quasi-perfect.
+The proof of ($\star$3) is similar to that of Lemma˜3.2; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work.
+Pick a point $p\in U$ and look at the fibered square
+[…]
+where $s_{p}$ is the natural morphism and $f_{p}$ and $t_{p}$ are the projections.
+Again, choose any compact generator $Q$ of $Y$.
+The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect.
+As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper.
+Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma˜2.1 and $f$ is quasi-perfect by Lemma˜3.1.
+∎
+Corollary 3.6.
+Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜3.4. Then the locus
+$S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
+is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
+Proof.
+By the proof of Theorem˜3.4 (i.e. using Lemma˜3.2) we know that
+$S_{X}$= $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
+$=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{h}\textrm{ is quasi-perfect}\right\}$
+and
+$S_{U}$= $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
+$=$ $\left\{u\in U\mid\text{base change to }\widehat{\mathcal{O}}_{U,u}\textrm{ is quasi-perfect}\right\}.$
+Let us start by showing that
+(3.3) $s^{-1}(S_{X})=S_{U}.$
+To this end, observe by Equation˜2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$.
+As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma˜3.2.
+Therefore,
+$s^{-1}(S_{X})$ $=\left\{u\in U\mid\text{base change to }\mathcal{O}_{X,s(u)}^{sh}\textrm{ is quasi-perfect}\right\}$
+$=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}^{sh}\textrm{ is quasi-perfect}\right\}$
+$=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}=S_{U}.$
+As a consequence of Equation˜3.3 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open.
+Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
+Therefore, we may and will simply assume that $X$ is an affine scheme from now on.
+Choose any compact generator $Q$ of $Y$.
+As $f$ is proper, $\mathbf{R}f_{\ast}Q$ is bounded coherent and so by [LET21, Proposition 3.5] it follows that
+$S^{\prime}:=\left\{x\in X\mid(\mathbf{R}f_{\ast}Q)_{x}\in\langle\mathcal{O}_{X,x}\rangle=\operatorname{Perf}(\mathcal{O}_{X,x})\right\}$
+is a Zariski open subset of $X$. Considering the fibered square
+[…]
+we see that, using flatness of $s_{x}$ and that it is affine so that $t_{x}^{\ast}Q$ is a compact generator for $Y\times_{X}\operatorname{Spec}(\mathcal{O}_{X,x})$,
+$x\in S\iff$ $\text{base change to }\mathcal{O}_{X,x}\textrm{ is quasi-perfect}$
+$\iff$ $\mathbf{R}(f_{x})_{\ast}t_{x}^{\ast}Q\text{ is perfect}$
+$\iff$ $(\mathbf{R}f_{\ast}Q)_{x}=s_{x}^{\ast}\mathbf{R}f_{\ast}Q\text{ is perfect}\iff x\in S^{\prime}$
+finishing the proof. ∎
+Remark 3.7.
+One can think of the subset $S$ appearing in Corollary˜3.6 as the ‘quasi-perfect locus of $f$’.
+It follows from the first part of the proof that in general, i.e.
+regardless of properness of $f$, one can compute it étale locally (Equation˜3.3) and that the specific choice of $A_{\#}$ does not matter.
+Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
+It is also worthwhile to point out that, quite generally for any morphisms $f\colon Y\to X$ (of quasi-compact quasi-separated algebraic spaces) with separated target, we have an inclusion of the regular locus $\operatorname{reg}(X)\subseteq S$.
+Indeed, if $p\in|X|$ is a regular point, its étale local ring is regular and so the base change to it is quasi-perfect.
+As an application of the ‘quasi-perfect locus’, we show it can be used to descent properties of the regular locus along proper morphisms.
+Moreover, note that the result requires no separations assumptions on the target and that the assumption on $X$ in the statement is a necessary condition due to the conclusion of the proposition.
+Proposition 3.8.
+Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [STA26, Tag 04NA] for the relevant definition.
+whose étale local ring is reduced (e.g.
+this is the case when $X$ is reduced).
+Suppose further that $f\colon Y\to X$ is a proper surjective morphism.
+Then the regular locus of $X$ contains a nonempty open subset (i.e.
+is $J\textrm{-}0$) whenever $Y$ has this property.
+Proof.
+The proof starts by a series of reductions. Let us denote by $\eta$ the codimension zero point whose étale local ring is reduced.
+  • As $X$ admits an open dense subscheme, which in particular contains $\eta$, see e.g.
+[STA26, Tag 06NH], we may assume that $X$ is a scheme.
+Moreover, by choosing any affine open neighborhood $U$ of $\eta$ and by restricting to $f^{-1}(U)\to U$, we may assume $X$ is affine and, in particular, separated.
+  • Next, let $S$ denote the open subset from Corollary˜3.6; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty.
+It follows from Theorem˜3.4 that the restriction $f^{-1}(S)\to S$ is quasi-perfect.
+So we may assume $f$ is quasi-perfect.
+  • Let $V$ be a non-empty open subset of $Y$ contained in the regular locus, which exists by assumption.
+If $V\neq Y$, then $Y\setminus V$ is a properly contained closed subset of $Y$.
+As $f$ is proper, $f(Y\setminus V)$ is closed in $X$.
+Moreover, $f^{-1}(X\setminus f(Y\setminus V))\subset V$; so restricting to the open subset $X\setminus f(Y\setminus V)\subset X$ allows us to reduce to the setting where $f$ is a quasi-perfect proper surjective morphism from a regular scheme.
+(Note that whether or not $\eta\in X\setminus f(Y\setminus V)$ does not matter, the existence of $\eta$ was needed in the previous bullet.)
+In order to finish the proof, is suffices to show that $X$ is regular.
+So, let $p\in X$ be a closed point and denote the associated closed immersion by $i\colon\mathrm{Spec}(\kappa(p))\to X$.
+Consider the fibered square
+[…]
+Clearly, $\kappa(p)$ being a field, $\mathbf{R}f^{\prime}_{\ast}\mathcal{O}_{Y\times_{X}}\operatorname{Spec}(\kappa(p))$ is a (non-zero, as $f$ is surjective!) finite direct sum of shifts of $\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$.
+Thus, commutativity of the diagram shows we have $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}\in\langle\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{1}$.
+Regularity of $Y$ ensures that $\operatorname{Perf}(Y)=D^{b}_{\operatorname{coh}}(Y)$ and quasi-perfectness of $f$ tells us $\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$.
+Consequently, $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}\in\operatorname{Perf}(X)$, which ensures that $p$ is a regular point of $X$.
+As we have shown every closed point of $X$ is regular, it follows that $X$ is regular as desired.
+∎
+Appendix A Quasi-perfect blowups detect regularity
+The goal of this appendix is to prove Proposition˜A.3, which gives a relationship between regularity of a Noetherian algebraic space $X$ and having ‘enough’ quasi-perfectness of morphisms with target $X$.
+In an earlier version of our work, we recorded Proposition˜A.3 as an observation.
+However, Bhatt privately communicated to us preliminary work of his, which is to appear later, that subsumes Proposition˜A.3.
+However, we felt it worthwhile to record this (and some of the lemmas used in its proof) in an appendix.
+We start with the following lemma, which should be compared to [STA26, Tag 00OC] and gives a ‘spacy’ version of this.
+Lemma A.1.
 Let $X$ be a decent (e.g.
@@
 ∎
-Lemma 3.2.
-Let $X$ be Noetherian algebraic space $X$. If the blowup of $X$ along a closed point $p\in|X|$ induces a quasi-perfect morphisms, then $p$ is regular.
+The following lemma is the main crucial, albeit simple, observation.
+Lemma A.2.
+Let $X$ be Noetherian algebraic space $X$. If the blowup of $X$ along a closed point $p\in|X|$ induces a quasi-perfect morphism, then $p$ is regular.
 Proof.
@@
 where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor.
-By definition of blowing up the exceptional divisor $E$ is an effective Cartier divisor, and so it follows from the distinguished triangle
+By definition of blowing up, the exceptional divisor $E$ is an effective Cartier divisor, and so it follows from the distinguished triangle
 $\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to\mathcal{O}_{X^{\prime}}(-E)[1].$
@@
 Hence, as $f$ is quasi-perfect by assumption, $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$ is perfect too.
-To conclude, observe from the commutativity of Diagram˜3.1 that
+To conclude, observe from the commutativity of Diagram˜A.1 that
 $0$ $\not=\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$
@@
 Hence $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect.
-Thus, $p$ is regular by Lemma˜3.1, and we are done.
+Thus, $p$ is regular by Lemma˜A.1, and we are done.
 ∎
-Proposition 3.3.
+Proposition A.3.
 A Noetherian algebraic space $X$ is regular if and only if the blowup of $X$ along any closed point is a quasi-perfect morphism.
@@
 $\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\subseteq D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X).$
-Hence, any blowup is quasi-perfect as these are proper. The converse follows from Lemma˜3.2 as it suffices to check regularity at closed points (see e.g. [DLR+25, Lemma 3.2]). ∎
-Lemma 3.4.
+Hence, any blowup is quasi-perfect as these are proper. The converse follows from Lemma˜A.2 as it suffices to check regularity at closed points (see e.g. [DLM+25, Lemma 3.2]). ∎
+Recall that a morphism of integral algebraic spaces $f\colon Y\to X$ is a modification (resp.
+alteration) if it is proper and birational, i.e.
+generically an isomorphism (resp.
+generically finite).
+See e.g.
+[STA26, Tag 0ACU] and [STA26, Tag 0AD7] for some characterizations.
+Corollary A.4.
+An integral Noetherian algebraic space $X$ is regular if and only if every alteration (resp. modification) of $X$ is quasi-perfect.
+Proof.
+This follows immediately from Lemma˜A.2.
+Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma˜A.5 below.
+Moreover, as modifications are alterations the case of alterations also follows.
+∎
+Lemma A.5.
 The blowup $f\colon X^{\prime}\to X$ of an integral algebraic space along a nonzero quasi-coherent ideal sheaf is integral. In particular, when the ideal is of finite type, it is a modification.
@@
 To see that $f$ is a modification when the ideal is of finite type, note that in this case $f$ is proper by [STA26, Tag 085Z]. ∎
-Corollary 3.5.
-An integral Noetherian algebraic space $X$ is regular if and only if every alteration (resp. modification) of $X$ is quasi-perfect.
-Proof.
-This follows immediately from Lemma˜3.2.
-Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma˜3.4.
-Moreover, as modifications are alterations the case of alterations also follows.
-∎
-Remark 3.6.
-In the special case where there exists an alteration or modification $f\colon Y\to X$ of algebraic spaces where $Y$ is regular and $f$ quasi-perfect, then one can show $X$ regular more easily using ‘proper descent’.
-For example, this follows from [HLL+25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity.
-So, the novelty of Corollary˜3.5 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
-4. On proper quasi-perfect morphisms
-Here we study quasi-perfect morphisms which are proper. We start with a few lemmas that allow one to check quasi-perfectness using flat covers.
-Lemma 4.1.
-A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if and only if $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
-Proof.
-If $f$ is quasi-perfect, it is clear that $\mathbf{R}f_{\ast}P\in\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$.
-For, the converse direction, let $P$ be compact generator on $Y$.
-Our hypothesis is $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$.
-The desired claim then follows from the following string of inclusions:
-$\mathbf{R}f_{\ast}\operatorname{Perf}(Y)=\mathbf{R}f_{\ast}\langle P\rangle\subseteq\langle\mathbf{R}f_{\ast}P\rangle\subseteq\operatorname{Perf}(X).\qed$
-Lemma 4.2.
-Let $f\colon Y\to X$ be a morphism of quasi-compact quasi-separated algebraic spaces and let $s\colon U\to X$ be a flat morphism (from an algebraic space). Consider the fibered square
-[…]
-  1. (1)
-    If $f$ is quasi-perfect and $s$ is quasi-affine, then $f^{\prime}$ is quasi-perfect.
-  2. (2)
-    If $s$ is surjective and $f^{\prime}$ is quasi-perfect, then $f$ is quasi-perfect.
-Proof.
-To show the first claim, first note that when $s$ is an quasi-affine morphism so is its base change $s^{\prime}$.
-Consequently, $\mathbf{R}s^{\prime}_{\ast}$ is conservative (see e.g.
-[HR17, Corollary 2.8]) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g.
-[STA26, Tag 0E4R]).
-Thus, as $f$ is assumed quasi-perfect, $\mathbf{R}f_{\ast}Q$ is perfect and so also is its pullback $s^{\ast}\mathbf{R}f_{\ast}Q$.
-From flat base change, it follows that
-(4.2) $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q\cong s^{\ast}\mathbf{R}f_{\ast}Q.$
-Hence, Lemma˜4.1 implies $f^{\prime}$ is a quasi-perfect morphism.
-For the second claim, assume now that $f$ is surjective and $f^{\prime}$ is quasi-perfect.
-Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation˜4.2.
-As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact.
-Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma˜2.1.
-By Lemma˜4.1, we have that $f$ is quasi-perfect, which completes the proof.
-∎
-The following is a variation on the previous.
-Lemma 4.3.
-Let $f\colon Y\to X$ be a morphism of quasi-compact quasi-separated algebraic spaces with $X$ an affine scheme.
-Suppose $s\colon U\to X$ is a flat surjective morphism from an affine scheme.
-Then $f$ is quasi-perfect if and only if the base change of $f$ along $s$ is quasi-perfect.
-Proof.
-The forward direction follows from Lemma˜4.2.
-For the converse, denote the base change of $f$ along $s$ as in Diagram˜4.1.
-Again, choose a compact generator $Q$ of $Y$.
-By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation˜4.2, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect.
-However, as $s$ corresponds to a faithfully flat ring map, it follows from [STA26, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect.
-Thus, Lemma˜4.1 shows that $f$ is quasi-perfect.
-∎
-Using Lemma˜4.2, we obtain an alternative proof of the fact that regularity can be checked fpqc locally. Before doing so, we recall a fact whose proof we add for lack of finding a reference.
-Lemma 4.4.
-Let $X$ be a quasi-separated algebraic space. Then any morphism from an affine scheme to $X$ is quasi-affine.
-Proof.
-Suppose $U\to X$ is a map from an affine scheme.
-So show it is affine, by definition, we need to show, for any affine scheme $V$ and morphisms $V\to X$, that $U\times_{X}V$ is quasi-affine.
-As the latter is just the pullback of the morphisms $U\times V\to X\times X$ along the diagonal $X\to X\times X$, this follows from the fact the diagonal of $X$ is quasi-affine, see e.g.
-[STA26, Tag 03HK & Tag 02LR].
-∎
-Corollary 4.5.
+To finish, using Lemmas˜3.2 and A.2, we give an alternative proof of the fact that regularity can be checked fpqc locally.
+Corollary A.6.
 Let $X$ be a Noetherian algebraic space. If there is an fpqc cover $\{f_{i}\colon U_{i}\to X\}_{i\in I}$ where each $U_{i}$ is regular, then $X$ is regular.
@@
 take étale surjective morphisms from affine schemes).
-Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is quasi-affine by Lemma˜4.4 as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme.
-To show $X$ is regular, it suffices by Lemma˜3.2 to show that blowups along nonzero ideal sheaves are quasi-perfect.
+Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is quasi-affine by Lemma˜3.5 as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme.
+To show $X$ is regular, it suffices by Lemma˜A.2 to show that blowups along nonzero ideal sheaves are quasi-perfect.
 However, as blowups commute with flat base change, see e.g.
@@
 Consequently, as $U$ is regular, it is quasi-perfect.
-Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma˜4.2, showing the claim.
+Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma˜3.2, showing the claim.
 ∎
-Remark 4.6.
+Remark A.7.
 In fact, when looking at fppf covers it is enough to only assume that $X$ is quasi-compact and decent in the previous by making use of some descent.
 Indeed, it follows from the existence of an fppf cover as in the statement that $X$ is locally Noetherian (as this is fppf local and regular algebraic spaces are always locally Noetherian) and hence Noetherian.
-There are three important cases of faithfully flat morphisms that we will apply Lemma˜4.3.
-Example 4.7.
-Let $(R,\mathfrak{m})$ be a Noetherian local ring. To this one can associate a few other Noetherian local rings:
-  1. (1)
-    the $\mathfrak{m}$-adic completion of $R$, denoted $\widehat{R}$,
-  2. (2)
-    the Henselization of $R$, denoted $R^{h}$, and
-  3. (3)
-    the strict Henselization of $R$, denoted $R^{sh}$.
-There are local ring morphisms $R\to\widehat{R}$ and $R\to R^{h}\to R^{sh}$ which are flat and consequently faithfully flat; see e.g.
-[STA26, Tag 00MC and Tag 07QM].
-(As an aside, there is also a flat local ring morphism $R^{h}\to\widehat{R}$, but we do not need it).
-We now move to our next main result.
-Theorem 4.8.
-Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
-  1. (1)
-    $f$ is quasi-perfect,
-  2. (2)
-    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in|X|$.
-Furthermore, if $s\colon U\to X$ is an étale surjective morphism from a quasi-compact scheme, one can take $A_{\#}\in\{\mathcal{O}_{U,\#},\widehat{\mathcal{O}}_{U,\#}\}$ and let $p$ run over all points in $U$.
-Proof.
-[…]
-[…]
-($\star$) If the base change of $f$ to $\mathcal{O}_{U,p}$ is quasi-perfect for all $p\in U$, then $f$ is quasi-perfect.
-The proof of ($\star$4) is similar to that of Lemma˜4.2; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work.
-Pick a point $p\in U$ and look at the fibered square
-[…]
-where $s_{p}$ is the natural morphism and $f_{p}$ and $t_{p}$ are the projections.
-Again, choose any compact generator $Q$ of $Y$.
-The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect.
-As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper.
-Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma˜2.1 and $f$ is quasi-perfect by Lemma˜4.1.
-∎
-Corollary 4.9.
-Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
-$S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
-is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
-Proof.
-By the proof of Theorem˜4.8 (i.e. using Lemma˜4.3) we know that
-$S_{X}$= $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
-$=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{h}\textrm{ is quasi-perfect}\right\}$
-and
-$S_{U}$= $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
-$=$ $\left\{u\in U\mid\text{base change to }\widehat{\mathcal{O}}_{U,u}\textrm{ is quasi-perfect}\right\}.$
-Let us start by showing that
-(4.3) $s^{-1}(S_{X})=S_{U}.$
-To this end, observe by Equation˜2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$.
-As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma˜4.3.
-Therefore,
-$s^{-1}(S_{X})$ $=\left\{u\in U\mid\text{base change to }\mathcal{O}_{X,s(u)}^{sh}\textrm{ is quasi-perfect}\right\}$
-$=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}^{sh}\textrm{ is quasi-perfect}\right\}$
-$=\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}=S_{U}.$
-As a consequence of Equation˜4.3 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open.
-Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
-Therefore, we may and will simply assume that $X$ is an affine scheme from now on.
-Choose any compact generator $Q$ of $Y$.
-As $f$ is proper, $\mathbf{R}f_{\ast}Q$ is bounded coherent and so by [LET21, Proposition 3.5] it follows that
-$S^{\prime}:=\left\{x\in X\mid(\mathbf{R}f_{\ast}Q)_{x}\in\langle\mathcal{O}_{X,x}\rangle=\operatorname{Perf}(\mathcal{O}_{X,x})\right\}$
-is a Zariski open subset of $X$. Considering the fibered square
-[…]
-we see that, using flatness of $s_{x}$ and that it is affine so that $t_{x}^{\ast}Q$ is a compact generator for $Y\times_{X}\operatorname{Spec}(\mathcal{O}_{X,x})$,
-$x\in S\iff$ $\text{base change to }\mathcal{O}_{X,x}\textrm{ is quasi-perfect}$
-$\iff$ $\mathbf{R}(f_{x})_{\ast}t_{x}^{\ast}Q\text{ is perfect}$
-$\iff$ $(\mathbf{R}f_{\ast}Q)_{x}=s_{x}^{\ast}\mathbf{R}f_{\ast}Q\text{ is perfect}\iff x\in S^{\prime}$
-finishing the proof. ∎
-Remark 4.10.
-One can think of the subset $S$ appearing in Corollary˜4.9 as the ‘quasi-perfect locus of $f$’.
-It follows from the first part of the proof that in general, i.e.
-regardless of properness of $f$, one can compute it étale locally (Equation˜4.3) and that the specific choice of $A_{\#}$ does not matter.
-Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
-It is also worthwhile to point out that, quite generally for any morphisms $f\colon Y\to X$ (of quasi-compact quasi-separated algebraic spaces) with separated target, we have an inclusion of the regular locus $\operatorname{reg}(X)\subseteq S$.
-Indeed, if $p\in|X|$ is a regular point, its étale local ring is regular and so the base change to it is quasi-perfect.
-As an application of the ‘quasi-perfect locus’, we show it can be used to descent properties of the regular locus along proper morphisms.
-Moreover, note that the result requires no separations assumptions on the target and that the assumption on $X$ in the statement is a necessary condition due to the conclusion of the proposition.
-Proposition 4.11.
-Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [STA26, Tag 04NA] for the relevant definition.
-whose étale local ring is reduced (e.g.
-this is the case when $X$ is reduced).
-Suppose further that $f\colon Y\to X$ is a proper surjective morphism.
-Then the regular locus of $X$ contains a nonempty open subset (i.e.
-is JJ-0) whenever $Y$ has this property.
-Proof.
-The proof starts by a series of reductions. Let us denote by $\eta$ the codimension zero point whose étale local ring is reduced.
-  • As $X$ admits an open dense subscheme, which in particular contains $\eta$, see e.g.
-[STA26, Tag 06NH], we may assume that $X$ is a scheme.
-Moreover, by choosing any affine open neighborhood $U$ of $\eta$ and by restricting to $f^{-1}(U)\to U$, we may assume $X$ is affine and, in particular, separated.
-  • Next, let $S$ denote the open subset from Corollary˜4.9; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty.
-It follows from Theorem˜4.8 that the restriction $f^{-1}(S)\to S$ is quasi-perfect.
-So we may assume $f$ is quasi-perfect.
-  • Let $V$ be a non-empty open subset of $Y$ contained in the regular locus, which exists by assumption.
-If $V\neq Y$, then $Y\setminus V$ is a properly contained closed subset of $Y$.
-As $f$ is proper, $f(Y\setminus V)$ is closed in $X$.
-Moreover, $f^{-1}(X\setminus f(Y\setminus V))\subset V$; so restricting to the open subset $X\setminus f(Y\setminus V)\subset X$ allows us to reduce to the setting where $f$ is a quasi-perfect proper surjective morphism from a regular scheme.
-(Note that whether or not $\eta\in X\setminus f(Y\setminus V)$ does not matter, the existence of $\eta$ was needed in the previous bullet.)
-In order to finish the proof, is suffices to show that $X$ is regular.
-This follows from ‘proper descent’, as already indicated in Remark˜3.6; but we give a self-contained proof, more in line with that of Lemma˜3.2, for the convenience of the reader.
-So, let $p\in X$ be a closed point and denote the associated closed immersion by $i\colon\mathrm{Spec}(\kappa(p))\to X$. Consider the fibered square
-[…]
-Clearly, $\kappa(p)$ being a field, $\mathbf{R}f^{\prime}_{\ast}\mathcal{O}_{Y\times_{X}}\operatorname{Spec}(\kappa(p))$ is a (non-zero, as $f$ is surjective!) finite direct sum of shifts of $\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$.
-Thus, commutativity of the diagram shows we have $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}\in\langle\mathbf{R}f_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{1}$.
-Regularity of $Y$ ensures that $\operatorname{Perf}(Y)=D^{b}_{\operatorname{coh}}(Y)$ and quasi-perfectness of $f$ tells us $\mathbf{R}f_{\ast}\operatorname{Perf}(Y)\subseteq\operatorname{Perf}(X)$.
-Consequently, $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}\in\operatorname{Perf}(X)$, which ensures that $p$ is a regular point of $X$.
-As we have shown every closed point of $X$ is regular, it follows that $X$ is regular as desired.
-∎